### v1

# On the Approximation Power of SiLU Networks: Exponential Rates and Depth Efficiency

Abstract
This article establishes a comprehensive theoretical framework demonstrating that SiLU (Sigmoid Linear Unit) activation networks achieve exponential approximation rates for smooth functions with explicit and improved complexity control compared to classical ReLU-based constructions. We develop a novel hierarchical construction beginning with an efficient approximation of the square function $x^{2}$ more compact in depth and size than comparable ReLU realizations, such as those given by Yarotsky. This construction yields an approximation error decaying as $\mathcal{O}(\omega^{-2k})$ using networks of depth $\mathcal{O}(1)$. We then extend this approach through functional composition to establish sharp approximation bounds for deep SiLU networks in approximating Sobolev-class functions, with total depth $\mathcal{O}(1)$ and size $\mathcal{O}(\varepsilon^{-d/n})$.
Keywords: Neural network approximation, SiLU activation, Sobolev spaces, Complexity bounds.
1 Introduction
The study of approximation capabilities of neural networks is a central topic in the theory of deep learning. Since the classical results of Cybenko [cybenko1989approximation], Hornik [hornik1991approximation], and Funahashi [funahashi1989approximate], it is known that single-hidden-layer neural networks (NN) with nonlinear activation functions can approximate any continuous function on a compact domain to arbitrary accuracy. These universal approximation theorems, however, do not address efficiency, i.e., how network size relates to approximation error for specific function classes.
Barron [cheang2001penalized] introduced risk bounds for function estimation with single-hidden-layer networks, using a penalized least-squares criterion for model selection. The development of deep networks and the ReLU activation function later led to sharper approximation results. Yarotsky [yarotsky2017error] established exponential expressivity of ReLU networks for smooth functions, and Petersen and Voigtlaender [petersen2018optimal] derived optimal bounds for function classes with variable regularity. The roles of network depth and width in approximation efficiency were clarified in [telgarsky2016benefits, lu2017expressive, arora2016understanding].
In addition to ReLU, other perspectives on neural network expressivity have been developed. Daniely, in [daniely2020neural], analyzed learning mechanisms beyond over-parameterization. More recently, smooth activation functions have been reconsidered, after initially being set aside, due to their empirical performance. The SiLU (or swish) activation [elfwing2018sigmoid, ramachandran2017searching] and its variant GELU [hendrycks2016gaussian, lee2023gelu] have been adopted in modern architectures for variational surrogate model [ayena:hal-05178445].
On the theoretical side, several recent works have examined the approximation and learning properties of smooth activations. Yet, a systematic characterization of the approximation power of SiLU networks, comparable to what is known for ReLU, has not been established.
The purpose of this work is to address this gap by providing a rigorous analysis of the approximation properties of SiLU-activated networks, including convergence rates and complexity bounds, thereby situating them within the general framework of neural network approximation theory. Our work contributes to the growing literature on the expressive power of neural networks [petersen2018optimal, lu2017expressive, telgarsky2016benefits] by providing explicit construction methods with guaranteed error bounds. The results have implications for understanding why shallow or deep networks with smooth activations perform so effectively in practice and provide theoretical guidance for the design of network architectures. The constructive nature of our proofs offers potential pathways for developing specialized networks for scientific computing applications where high precision is required.
The remainder of this article is organized as follows. Section 2 provides essential background on neural network definitions and uniform approximation concepts, establishing the foundational framework for our analysis. In Section 3, we present our main contributions, beginning with the fundamental approximation of the square function using SiLU networks, which serves as the building block for subsequent results. We then develop efficient approximations for multiplication operations, arbitrary monomials, and general polynomials, demonstrating how these constructions naturally lead to recurrent neural network architectures with shared parameters. This section concludes with our main universal approximation theorems for continuous functions and for functions defined in the unit ball of Sobolev spaces, providing explicit complexity bounds. Section 4 discusses practical implications and presents experimental results validating our theoretical findings comparatively to ReLU networks.
2 Definition of neural network and uniform approximation
A neural network
A neural network of depth $m$ (i.e., $m$ layers including input and output, or (m1)(m-1) hidden layers) is a function $N:\mathbb{R}^{p}\to\mathbb{R}^{d}$ defined as the composition of affine maps and nonlinear activation functions. More precisely, the network is given by
$N(\mathbf{x})=(\phi_{m}\circ h_{m})\circ(\phi_{m-1}\circ h_{m-1})\circ\cdots\circ(\phi_{1}\circ h_{1})(\mathbf{x}),$
where for each $i\in\{1,\dots,m\}$,
  • the $i$-th hidden layer contains $n_{i}$ neurons, corresponding to the $n_{i}$ coordinates of $\phi_{i}\circ h_{i}(\mathbf{x}_{i})$ computation is given by an affine transformation $h_{i}:\mathbb{R}^{n_{i-1}}\to\mathbb{R}^{n_{i}}$ of the form
    $h_{i}(\mathbf{x}_{i})=W_{i}\mathbf{x}_{i}+b_{i},$
[…]
  • and the dimensions satisfy $n_{0}=p$ and $n_{m}=d$.
Let $f:[-B,B]^{p}\to\mathbb{R}$ be a function. We say that a function $g$ approximates $f$ with precision $\epsilon>0$ if
$\sup_{\mathbf{x}\in[-B,B]^{d}}|g(\mathbf{x})-f(\mathbf{x})|\leq\epsilon$
This inequality ensures that the approximation error is uniformly bounded by $\epsilon$ over the entire domain $B,B]^{d}$, thereby guaranteeing pointwise accuracy throughout the considered region. In the context of neural networks, this means that the network $g$ must faithfully capture the behavior of $f$ over the entire domain without exception. The complexity of such network (its depth and size) will depend on the regularity of $f$ and the required precision $\epsilon$.
Convention. If $\mathbf{x}=(x_{1},\dots,x_{p})\in\mathbb{R}^{p}$, then
$\text{SiLU}(\mathbf{x}):=\big(\text{SiLU}(x_{1}),\dots,\text{SiLU}(x_{p})\big)\in\mathbb{R}^{p},$
that is, SiLU is applied componentwise to vectors, where
$\forall\ x\in\mathbb{R},\quad\text{SiLU}(x)=\frac{x}{1+e^{-x}}.$
Figure 1: Approximation of the function $\log(7+x)\cos(x^{3})$ by ReLU neural networks (left, Mse =1.69) and SiLU neural networks (right, MSE=1.63).
2.1 Motivations
The choice of activation function is a critical factor in the ability of neural networks to approximate complex functional behaviors. A comparison between two identical architectures ($12$ hidden layers and one linear output layer with 64 neurons per hidden layer) using the SiLU activation function and the ReLU activation function ($\max(0,x)$), respectively, reveals experimentally that the SiLU version provides superior approximation accuracy for equal complexity.
To illustrate this, consider the function
$T_{1}(x)=\log(7+x)\cdot\cos(x^{3}),$
defined on a compact interval. The network using SiLU produces a lower approximation error and yields a smoother curve, more closely matching the local variations of $f$. This superiority is explained by several fundamental properties of the SiLU function: its regularity ($C^{\infty}$), its differentiability over its entire domain, and a better-behaved gradient specifically, non-zero over $\mathbb{R}$, unlike ReLU.
These observations motivate the design of compact and optimized neural networks that specifically leverage SiLU activation for function approximation. Such an approach could be to:
  • Reduce the size of networks required to achieve a given precision,
  • Improve generalization capabilities and numerical stability,
  • Achieve enhanced performance on highly regular functions.
3 Our theoritical results
3.1 SiLU is robust for polynomial approximation
The following theorem establishes the ability of SiLU neural networks to efficiently approximate the square function, which serves as a fundamental building block for approximating arbitrary polynomials. This result demonstrates the robustness of SiLU activation for polynomial approximation, as it achieves exponential convergence with constant-depth networks.
Theorem 1.
For any $B>0$ and target accuracy $\epsilon\in(0,1)$ , there exist constants $C_{1}>0$, $\omega_{1}>1$ and a family of feedforward SiLU neural networks $\{Q_{k}\}_{k\in\mathbb{N}}$, with one hidden layer with $3$ neurons each, such that for $k\geq\max\left\{\left\lceil\frac{\ln C_{1}-\ln\epsilon}{2\ln\omega_{1}}\right\rceil,\left\lceil\frac{\ln 2B}{2\ln\omega_{1}}\right\rceil\right\}$
$\sup_{|x|\leq B}|Q_{k}(x)-x^{2}|\leq\epsilon,$
where $\lceil\cdot\rceil$ is the ceiling function. Moreover, the network $Q_{k}$ has depth $\mathcal{O}(1)$ and size $\mathcal{O}(1)$.
Proof.
The complete proof is provided in Appendix A.1. We then construct the function
$g_{a}(x)=(x+a)\sigma(x+a)+(-x+a)\sigma(-x+a)-2a\sigma(a),$
which eliminates the odd-order terms, leaving only even-order terms starting from $x^{2}$. The dominant coefficient of $x^{2}$ is $K(a)=2\sigma^{\prime}(a)+a\sigma^{\prime\prime}(a)$, which is non-zero for an appropriate choice of $a$.
We normalize $g_{a}(x)$ to isolate the $x^{2}$ term and express the residual error as a series of higher even powers. To reduce this error, we introduce a sequence of functions $Q_{k}$ defined recursively via a scaling transformation:
$Q_{k}(x)=\beta^{-2k}f_{a}(\beta^{k}x),$
with $\beta<1$. This construction exponentially compresses the error terms, as the coefficients of higher-order terms are multiplied by $\beta^{k(j-2)}$ for the term in $x^{j}$.
Over any interval [B,B][-B,B], the error $|Q_{k}(x)-x^{2}|$ is bounded by a sum of terms that decay exponentially with $k$, ensuring uniform convergence with an error of $\mathcal{O}(\beta^{2k})$. Finally, we show that the function $Q_{k}$ can be implemented by a SiLU neural network of constant depth and size, using a linear combination of three SiLU evaluations with carefully chosen parameters
$Q_{k}(x)=\gamma\cdot SiLU(Wx+b)$ (1)
with parameters such that the weights $W=[\beta^{k},-\beta^{k},0]^{\top}$, $\gamma=\frac{\beta^{-2k}}{K(a)}[1,1,-2]$, and the bias $b=[a,a,a]^{\top}$. ∎
Figure 2: Architecture of NN $Q_{k}$ (Theorem 1).
Figure 3: Approximation of the function $x^{2}$ by the functions $Q_{k}$ (Theorem 1).
We use the classical algebraic identity
$xy=\frac{1}{4}[(x+y)^{2}-(x-y)^{2}]$
to establish the ability of neural networks with SiLU activation functions to efficiently approximate the multiplication of two variables. It demonstrates that the approximation error decreases exponentially fast as the complexity of the network increases, all while using an architecture whose size and depth remain constant, independent of the required level of precision.
Theorem 2.
For every $B>0$ and target accuracy $\epsilon\in(0,1)$, there exist constants $C_{2}>0$, $\omega_{2}>1$ and a family of feedforward SiLU neural networks $\{M_{k}\}_{k\in\mathbb{N}}$, with at most one hidden layer with $4$ neurons each, such that for $k\geq\max\left\{\left\lceil\frac{\ln C_{2}-\ln\epsilon}{2\ln\omega_{2}}\right\rceil,\left\lceil\frac{\ln 2B}{2\ln\omega_{2}}\right\rceil\right\}$
$\sup_{|x|,|y|\leq B}|M_{k}(x,y)-xy|\leq\epsilon$
The network $M_{k}$ has depth $\mathcal{O}(1)$ and size $\mathcal{O}(1)$.
Proof.
The complete and detailed proof can be found in Appendix A.2. The proof is built upon the classical algebraic identity that decomposes the product into a difference of squares: $xy=\frac{1}{4}[(x+y)^{2}-(x-y)^{2}]$. The neural network $M_{k}$ is explicitly constructed to implement this identity using approximate squaring units.
The construction proceeds in three conceptual steps. First, the input pair $(x,y)$ is transformed linearly into the sum $(x+y)$ and the difference (xy)(x-y). Each of these two terms is then processed by a separate SiLU network, $Q_{k}$, which approximates the square function with an exponentially decaying error of $\mathcal{O}(\omega^{-2k})$, as per Theorem 1. Finally, the outputs of these two subnetworks are combined linearly to produce the final approximation of the product. This construction leads to an explicit formula for the entire network $M_{k}$, which can be expressed as a single, shallow SiLU network with a specific weight structure
$M_{k}(x,y)=\gamma\cdot\text{SiLU}(Wz+b)$ (2)
where $z=[x,y]^{\top}$ is the input vector. The parameters are meticulously chosen to encode the entire approximation process. The weight matrix
$W=\beta^{k}\begin{bmatrix}+1&+1\\ -1&-1\\ +1&-1\\ -1&+1\end{bmatrix}$
computes the four required linear combinations $(x+y)$, (x+y)-(x+y), (xy)(x-y), and (xy)-(x-y). The bias vector $b=[a,a,a,a]^{\top}$ is a constant. The output weight vector
$\gamma=\frac{\beta^{-2k}}{4K(a)}\begin{bmatrix}+1&+1&-1&-1\end{bmatrix}^{\top}$
performs the final linear combination, effectively calculating $\frac{1}{4}[Q_{k}(x+y)-Q_{k}(x-y)]$.
The exponential error rate for the product, $C_{2}\cdot\omega_{2}^{-2k}$, is a direct consequence of the error bounds from the two $Q_{k}$ subnetworks. Crucially, the entire computation is encapsulated in a network with constant depth, $\mathcal{O}(1)$, and constant size, $\mathcal{O}(1)$, as the two squaring networks are computed in parallel within this single layer. ∎
Figure 4: Architecture of NN $M_{k}$ (Theorem 2).
Figure 5: Approximation $M_{k}\left(\cos(x),\sin(x)\log(x^{2})\right)$ of the product of the functions $\cos(x)$ and $\sin(x)\log(x^{2})$ by the NN $\{M_{k}\}$ (Theorem 2).
This next result generalizes the previous approximation theorem from the product of two variables to arbitrary monomials $x^{m}$. It shows that SiLU networks can approximate any power function xmxm with exponential accuracy, using a network whose depth grows linearly with the degree $m$
Theorem 3.
For every $B>0,m\geq 3,$ and target accuracy $\epsilon\in(0,1)$, there exist constants $C_{m}>0$, $\omega_{m}>1$ and a family of feedforward SiLU neural networks $\{P_{m,k}\}_{k\in\mathbb{N}}$, with at most $2m$ hidden layers with $4$ neurons each, such that for $k\geq\max\left\{\left\lceil\frac{\ln C_{m}-\ln\epsilon}{2\ln\omega_{m}}\right\rceil,\left\lceil\frac{\ln 2B}{2\ln\omega_{m}}\right\rceil\right\}$
$\sup_{|x|\leq B}|P_{m,k}(x)-x^{m}|\leq\epsilon$
The network $P_{m,k}$ has depth $\mathcal{O}(m)$ and size $\mathcal{O}(1)$.
Proof.
This theorem is proven by induction on the degree $m$, building upon the multiplier network $M_{k}$ from the previous theorem. The base cases for $m=0,1,2$ are constructed directly (as a constant, the identity function, and the squaring network $Q_{k}$, respectively). The inductive step approximates $x^{m+1}$ by using the network $M_{k}$ to multiply the approximation of $x^{m}$ by $x$.
The resulting network $P_{m,k}$ for $m\geq 2$ has a recursive structure. It can be expressed as a composition of networks where each stage $i$ computes a product, leading to the final explicit architecture:
$P_{m,k}(x)=\gamma_{m}\cdot\text{SiLU}\left(W_{m}z_{m}+b_{m}\right)$ (3)
with the parameters defined as:
$W_{i}=\beta^{k}\begin{bmatrix}+1&+1\\ -1&-1\\ +1&-1\\ -1&+1\end{bmatrix},\quad\gamma_{i}=\frac{\beta^{-2k}}{4K(a)}\begin{bmatrix}+1\\ +1\\ -1\\ -1\end{bmatrix}^{\top},\quad b_{i}=[a,a,a,a]^{\top}.$ (4)
The input vector is $z_{i}=[x,\gamma_{i-1}\cdot\text{SiLU}(Wz_{i-1}+b_{i-1})]^{\top},\ \forall\ i\in\{2,3,\cdots,m\}$ such that $z_{1}=[x,1]^{\top}.$ This architecture ensures the network maintains a depth of $\mathcal{O}(m)$ and a size of $\mathcal{O}(1)$. The detailed proof, can be found in Appendix A.3
Figure 6: Approximation of the function $x^{7}$ by the functions $P_{m,k}$ (Theorem 3).
3.1.1 A Natural Recurrent Neural Network (RNN) Construction
A crucial observation emerges from our construction: approximation network parameters are naturally shareable. Consequently,
Corollary 1.
The resulting network $P_{m,k}$ of Theorem (3) is a shallow Recurrent Neural Network (RNN) with at most $2m$ hidden layer of $4$ neurons per step $i$ whose parameters are fixed.
Proof.
The parameters remain unchanged $W_{i}\equiv W,\;b_{i}\equiv b,\;\gamma_{i}\equiv\gamma$ for all $i$, making the architecture recurrent: it is exactly a Recurrent Neural Network (RNN) where the same operator is applied $m$ times. Our specific architecture is:
$z_{i}=[x,\ y_{i-1}]^{\top}\quad\text{and}\quad y_{i}=\gamma\cdot\text{SiLU}(Wz_{i}+b)$
The equations of our specific RNN at step $i$ are:
$\mathbf{h}_{i}$ $=\text{SiLU}\left(\mathbf{W}_{h}\cdot x+\mathbf{U}_{h}\cdot y_{i-1}+\mathbf{b}_{h}\right)$
$y_{i}$ $=\mathbf{W}_{y}\cdot\mathbf{h}_{i}$ (5)
With the constrained parameters:
$\mathbf{W}_{h}=\beta^{k}\begin{bmatrix}+1\\ -1\\ +1\\ -1\end{bmatrix},\ \mathbf{U}_{h}=\beta^{k}\begin{bmatrix}+1\\ -1\\ -1\\ +1\end{bmatrix},\ \mathbf{b}_{h}=[a,a,a,a]^{\top},\ \mathbf{W}_{y}=\frac{\beta^{-2k}}{4K(a)}\begin{bmatrix}+1&+1&-1&-1\end{bmatrix}$ (6)
And the initial states: $y_{0}=1$ (We can consider that the first operation $z_{1}=[x,1]^{\top}$ amounts to initializing the hidden state $h_{1}$ with $y_{0}=1$).
Figure 7: Approximation of the $x^{7}$ function by a recurrent neural network (5) with parameters given by (6)($k=3$).
$P_{m,k}$ is indeed a shallow RNN (a $2m$ hidden layer of $4$ neurons per step) whose parameters are fixed, shared, and not learned in this context, as they were chosen analytically to compute a specific function ($x^{m}$) rather than learned from data. ∎
This final result demonstrates that the ability to approximate efficiently monomials functions, as established in the previous theorem, directly extends to the approximation of arbitrary polynomials. The following theorem shows that a SiLU network can approximate any polynomial with exponential accuracy, where the complexity of the network scales linearly with the degree of the polynomial.
Theorem 4.
Let $N(x)=\sum_{m=0}^{M}a_{m}x^{m}$ be a polynomial of degree $M$. For every $B>0$ and target accuracy $\epsilon\in(0,1)$, there exist constants $C_{M}>0$, $\omega_{M}>1$ and a family of SiLU neural network $\{\mathcal{N}_{M,k}\}_{k\in\mathbb{N}}$, with at most $2M$ hidden layers with $4$ neurons each, such that for $k=\left\{\left\lceil\frac{\ln(C_{M}\cdot\max_{m}|a_{m}|)-\ln\epsilon}{2\ln\omega_{M}}\right\rceil,\left\lceil\frac{\ln 2B}{2\ln\omega_{M}}\right\rceil\right\}$
$\sup_{|x|\leq B}|\mathcal{N}_{M,k}(x)-N(x)|\leq\epsilon$
The network has depth $\mathcal{O}(M)$ and size $\mathcal{O}(1)$.
Proof.
This result is a direct consequence of the previous theorem on monomial approximation. The network $\mathcal{N}_{M,k}$ is constructed as a linear combination of the approximating networks $P_{m,k}$ for each monomial $x^{m}$, using the coefficients $a_{m}$ of the polynomial. The exponential error bound follows from the triangle inequality and the individual error bounds of each $P_{m,k}$. The depth of the network is dominated by the deepest monomial network, which is $O(M)$, while the size is $O(1)$ due to parameter sharing and the fact that the monomial networks $P_{M,k}$ that approximates monomial $x^{M}$ computes recursively approximated monomial of real monomial functions $x^{m},m\leq M.$ The complete proof, can be found in Appendix A.4.
Figure 8: Approximation of the function $x^{2}+x+1$ by the functions $\mathcal{N}_{M,k}$ (Theorem 4).
Remark 1.
In the case where the parameters $c_{i}$ are to be learned, if we have a dataset $(x_{j},y_{j}),j=1,2,\cdots,n$ such that $y_{j}=N(x_{j})$, then we can approximate the polynomial $y_{j}$ at $x_{j}$ by:
$\hat{y}_{j}=\begin{bmatrix}P_{0,k}(x_{j})&P_{1,k}(x_{j})&\cdots&P_{M,k}(x_{j})\end{bmatrix}\mathbf{c},$
with $\mathbf{c}^{\top}=[c_{0}\ c_{1}\ \ldots,\ c_{M}].$ The predicted $\mathbf{\hat{y}}\in\mathbb{R}^{n}$ output over the entire dataset can be written as a matrix product:
$\mathbf{\hat{y}}=\mathbf{P}\mathbf{c},$
where
$\mathbf{P}=\begin{bmatrix}P_{0,k}(x_{1})&P_{1,k}(x_{1})&\cdots&P_{M,k}(x_{1})\\ P_{0,k}(x_{2})&P_{1,k}(x_{2})&\cdots&P_{M,k}(x_{2})\\ \vdots&\vdots&\ddots&\vdots\\ P_{0,k}(x_{n})&P_{1,k}(x_{n})&\cdots&P_{M,k}(x_{n})\end{bmatrix}\in\mathbb{R}^{n\times(M+1)}.$
and $P_{0,k}(x_{j})=1$, for all $j\in\{1,2,\cdots,n\}$.
Obtaining the optimal parameters is subject to minimizing the mean squared error:
$\frac{1}{n}\|\mathbf{\hat{y}}-\mathbf{P}\mathbf{c}\|_{2}.$
Gradient descent can be applied to learn $\mathbf{c}$. If the matrix $\mathbf{P}^{\top}\mathbf{P}$ is invertible, the exact solution can be obtained in one step:
$\mathbf{c^{*}}=(\mathbf{P}^{\top}\mathbf{P})^{-1}\mathbf{P}^{\top}\mathbf{y}.$
Corollary 2 (Universal Polynomial Approximation via SiLU Recurrent Networks).
Let $N(x)=\sum_{m=0}^{M}a_{m}x^{m}$ be any polynomial of degree $M$. There exists a parameter-sharing recurrent neural network with SiLU activation that computes:
$\mathcal{N}_{M,k}(x)=\sum_{m=0}^{M}c_{m}y_{m}$
where the hidden states $y_{m}$ are computed sequentially through the recurrence:
$y_{0}$ $=1$
$\mathbf{h}_{m}$ $=\mathrm{SiLU}\left(\mathbf{W}_{h}x+\mathbf{U}_{h}y_{m-1}+\mathbf{b}_{h}\right)$
$y_{m}$ $=\mathbf{W}_{y}\mathbf{h}_{m}\quad\text{for }m=1,2,\ldots,M$
with fixed shared parameters:
$\mathbf{W}_{h}=\beta^{k}\begin{bmatrix}+1\\ -1\\ +1\\ -1\end{bmatrix},\ \mathbf{U}_{h}=\beta^{k}\begin{bmatrix}+1\\ -1\\ -1\\ +1\end{bmatrix},\ \mathbf{b}_{h}=\begin{bmatrix}a\\ a\\ a\\ a\end{bmatrix},\ \mathbf{W}_{y}=\frac{\beta^{-2k}}{4K(a)}\begin{bmatrix}+1&+1&-1&-1\end{bmatrix}$
such that
$\sup_{|x|\leq B}|\mathcal{N}_{M,k}(x)-N(x)|\leq\epsilon.$
Least‑squares realisation. In the sense of the $\|\cdot\|_{2}$ norm, the optimal coefficients $\mathbf{c}^{\ast}=(c_{0}^{\ast},\ldots,c_{M}^{\ast})$ for approximating $N(x)$ from data $\{(x_{j},y_{j})\}_{j=1}^{n}$ with $y_{j}=N(x_{j})$ can be obtained in closed form. Provided the matrix $\mathbf{P}^{\top}\mathbf{P}$ is invertible, the solution is given by:
$\mathbf{c}^{\ast}=(\mathbf{P}^{\top}\mathbf{P})^{-1}\mathbf{P}^{\top}\mathbf{y},$
where the design matrix $\mathbf{P}\in\mathbb{R}^{n\times(M+1)}$ has entries $\mathbf{P}_{j,m}=y_{m}(x_{j})$. This approximation satisfies the error bound:
$\|\mathcal{N}_{M,k}(x)-N(x)\|_{2}\leq\epsilon.$
Proof.
The result follows directly from Theorem 4 by observing that the recurrent network’s sequential computation naturally generates the monomial basis $\{1,x,x^{2},\ldots,x^{M}\}$ through its hidden states $y_{m}$ (Figure 7). The closed-form solution for the coefficients is obtained by recognizing that the approximation problem reduces to linear regression in the feature space spanned by the network’s hidden states. The error bound is inherited from the monomial approximation bounds of the constituent networks. ∎
Remark 2.
This corollary reveals a remarkable property: a single recurrent network with fixed parameters can serve as a universal polynomial feature extractor. The network’s architecture effectively implements a polynomial basis expansion, where:
  • The depth ($M$ steps) determines the maximum polynomial degree
  • The hidden dimension (4 units) determines the approximation quality
  • The coefficients $\mathbf{c}$ can be learned efficiently via convex optimization
This provides a computationally efficient alternative to traditional polynomial regression while maintaining interpretability through the polynomial coefficients.
Building upon the previous theorem 3 that demonstrated the approximation of the monomial $x^{m}$ using a deep architecture with $m$ hidden layers of four neurons each, the current theorem represents a significant improvement by reducing the architecture to a single hidden layer of $m+1$ neurons. This structural optimization is achieved by systematically leveraging the compositional properties and finite difference calculus of the SiLU activation function, which allows the polynomial approximation to be condensed into a single non-linear transformation while maintaining exponential convergence. The shift from depth to a moderate increase in width offers a practical advantage for implementations where depth-related constraints, such as vanishing gradients or computational complexity, are more critical than the width of individual layers.
3.2 A single hidden layer SiLU Network can approximate $x^{m}$ with arbitrary precision on any compact set ?
Theorem 5.
For any $B>0$ and target accuracy $\epsilon\in(0,1)$ , there exist constants $C_{1}>0$, $\omega_{1}>1$ and a family of feedforward SiLU neural networks $\{Q^{m}_{k}\}_{k\in\mathbb{N}}$, with one hidden layer with $m+1$ neurons each, such that for $k=\max\left\{\left\lceil\frac{\ln mB-\ln 2}{\ln\omega}\right\rceil,\left\lceil\frac{ln\epsilon-\ln C_{1,m}}{\ln\omega}\right\rceil\right\}$
$\sup_{|x|\leq B}|Q^{m}_{k}(x)-x^{m}|\leq\epsilon,$
Moreover, the network $Q^{m}_{k}$ has depth $\mathcal{O}(1)$ and size $\mathcal{O}(m)$.
Proof.
The complete proof is given in Appendix A.5. The construction begins by defining a linear combination of SiLU activations, $g_{a,m}(x)$, whose Taylor expansion is analyzed using finite difference calculus (Appendix B). This reveals that $g_{a,m}(x)=K_{m}(a)x^{m}+\text{(higher-order terms)}$. The function $f_{a,m}(x)=g_{a,m}(x)/K_{m}(a)$ thus provides an $m$-th order approximation to $x^{m}$. An iterative refinement is then applied: defining $Q^{m}_{k}(x)=\beta^{-m}Q^{m}_{k-1}(\beta x)$ with $Q^{m}_{0}=f_{a,m}$. It is shown that this process exponentially suppresses the higher-order error terms. A detailed analysis of the Taylor coefficients and the application of Cauchy’s integral formula (Appendix C) yield an explicit bound on the approximation error, $\sup_{|x|\leq B}|Q^{m}_{k}(x)-x^{m}|\leq C_{1,m}\omega^{-k}$, proving that a target accuracy $\epsilon$ is achieved for a suitable parameter $k$. The network $Q^{m}_{k}$ is explicitly parameterized, confirming it is a shallow network with $\mathcal{O}(m)$ parameters. Then we have
$Q^{m}_{k}(x)=\gamma_{m}\cdot\text{SiLU}(W_{m}x+b_{m})$ (7)
with parameters such that the weights
$W=\beta^{k}\,[w_{0},w_{1},\dots,w_{m}]^{\top},\gamma=\frac{\beta^{-mk}}{K_{m}(a)}[c_{0},c_{1},\dots,c_{m}],$
and the bias
$b=a\mathbf{1}_{m+1}.$
with
$w_{j}=j-\frac{m}{2},\quad c_{j}=(-1)^{m-j}\binom{m}{j}$,
and where $\mathbf{1}_{m+1}$ is the column vector of size $m+1$ containing ones. ∎
Remark 3.
The condition $K_{m}(a)\neq 0$ draws our attention when $a=0.$ Indeed in this case,
$K_{m}(0)=$ $0\sigma^{(m)}(0)+m\sigma^{(m-1)}(0)$
$=m\sigma^{(m-1)}(0)$
$=mf^{(m-1)}(0),\quad\text{with}\quad f(x)=\sigma(x)-1/2$
Noting that f is an odd function, it follows that $f^{(m-1)}(-x)=(-1)^{m}f^{(m-1)}(x).$ Thus, if $m$ is odd and $m\geq 3$
$f^{(m-1)}(0)=(-1)f^{(m-1)}(0)\implies f^{(m-1)}(0)=\sigma^{(m-1)}(0)=K_{m}(0)=0.$
We have shown that a SiLU network can approximate uniformly on a compact set any monomial $x^{m}$, for an arbitrary degree $m$, with controlled depth and width. Since monomials form a natural generating family for spaces of regular functions through polynomial expansions, this polynomial-approximation capability will serve as the foundation for the analysis that follows. In the next section, we extend this elementary result to the fast approximation of continuous functions, and then to functions with higher regularity, by SiLU neural networks.
3.3 Fast approximation of continuous functions and general smooth functions by SiLU Neural Networks
We establishes the approximation capability of SiLU networks for continuous functions, quantifying the efficiency of this approximation. The following result demonstrates that any continuous function on a compact interval can be approximated with arbitrary precision $\varepsilon$ by a SiLU network, and it provides explicit bounds on the required depth and size of the network to achieve this error. The key insight is the combination of classical functions (like polynomial) approximation (Weierstrass theorem) with the efficient neural approximation of polynomials developed in the previous theorems. In this section, we approximate functions on a compact set using two approaches: the polynomial approach and the step function approach.
Let $K\subset\mathbb{R}$ be a compact set. We denote by $C(K)$ the space of continuous functions on $K$, equipped with the uniform norm
$\|f\|_{\infty}=\sup_{x\in K}|f(x)|.$
We define, for $\alpha\in\mathbb{R}$, and $\tau>0$, $\kappa>0$, the function:
$\phi^{+}_{\alpha,\kappa,\tau}(x)=w\bigg[\text{SiLU}\left(\kappa\left(1-2w\,\text{SiLU}\left(\frac{\kappa(\alpha-x)}{\tau}\right)\right)\right)\bigg]$
where $w=1/\kappa$.
Finally, for $N\in\mathbb{N}^{*}$, let
$\mathcal{F}_{N}=\bigg\{\sum_{i=1}^{N}c_{i}\,\phi_{\alpha_{i},\beta_{i},\tau,\kappa}(x)\ \bigg|\ c_{i},\alpha_{i},\beta_{i}\in\mathbb{R},\ \tau>0,\ \kappa>0,\beta_{i}>\alpha_{i}\bigg\},\ \text{and}\ \mathcal{F}=\bigcup_{N\geq 1}\mathcal{F}_{N},$
such that
$\phi_{\alpha,\beta,\tau,\kappa}(x)=\phi^{+}_{\alpha,\tau,\kappa}(x)-\phi^{+}_{\beta,\tau,\kappa}(x)$
We have the following theorem:
Theorem 6.
The set $\mathcal{F}$ is dense in $\big(C(K),\,\|\cdot\|_{\infty}\big)$.
In order to prove this theorem, let us take a closer look at these two propositions.
Proposition 1.
Let $a<b$. For every $\delta>0$, there exists $\tau>0$, $\kappa_{0}>0$ such that for all $\tau\in(0,\tau_{0})$, $\kappa\geq\kappa_{0}$,
$\sup_{x\in\mathbb{R}}\left|\phi_{a,b,\tau,\kappa}(x)-\mathbf{1}_{[a,b]}(x)\right|<\delta.$
$\phi_{a,b,\tau,\kappa}(x)$ is a SiLU-activated neural network, with $2$ hidden layers, width $2$ with total number of parameters $12$.
Proof.
For all $x\in\mathbb{R}$, we have this limit
$\lim_{\kappa\to\infty}\frac{\text{SiLU}(\kappa x)}{\kappa}=\max(0,x)$ (8)
and the convergence is uniform on every compact subset of $\mathbb{R}$. We now give a simple proof of this fact.
Recall that
$f_{\kappa}(x)=\frac{\mathrm{SiLU}(\kappa x)}{\kappa}=x\,\sigma(\kappa x),\qquad(x)_{+}=\max(0,x).$
We treat separately the cases $x\geq 0$ and $x\leq 0$ to show that for every $R>0$,
$\sup_{|x|\leq R}\bigl|f_{\kappa}(x)-g(x)\bigr|\xrightarrow[\kappa\to\infty]{}0.$
Case $x\geq 0$.
In this region,
$|f_{\kappa}(x)-(x)_{+}|=|x\,(1-\sigma(\kappa x))|.$
Let $y=\kappa x$, so that $y\geq 0$ and $x=y/\kappa$. Then
$x(1-\sigma(\kappa x))=\frac{1}{\kappa}\,y\bigl(1-\sigma(y)\bigr).$
For $y\geq 0$ we have
$1-\sigma(y)=\frac{1}{1+e^{y}}\leq e^{-y},$
hence
$y\bigl(1-\sigma(y)\bigr)\leq ye^{-y}.$
The function $y\mapsto ye^{-y}$ is bounded on $0,\infty)$; it attains its maximum at $y=1$ with value $1/e$. Thus there is a constant $M\leq 1/e$ such that
$y\bigl(1-\sigma(y)\bigr)\leq M\quad\text{for all }y\geq 0.$
Therefore, for all $x\geq 0$,
$|f_{\kappa}(x)-(x)_{+}|\leq\frac{M}{\kappa}.$
Case $x\leq 0$.
Here, $(x)_{+}=0$ and
$|f_{\kappa}(x)-(x)_{+}|=|x|\,\sigma(\kappa x).$
Let $y=-\kappa x\geq 0$. Then
$|x|\,\sigma(\kappa x)=\frac{y}{\kappa}\,\sigma(-y)=\frac{1}{\kappa}\,\frac{y}{1+e^{y}}\leq\frac{1}{\kappa}\,ye^{-y}\leq\frac{M}{\kappa}.$
For every $R>0$ and all $x$ with $|x|\leq R$, we have
$|f_{\kappa}(x)-(x)_{+}|\leq\frac{M}{\kappa},$
hence
$\sup_{|x|\leq R}|f_{\kappa}(x)-(x)_{+}|\leq\frac{M}{\kappa}\xrightarrow[\kappa\to\infty]{}0.$
Thus the convergence is uniform on every compact set. Using the explicit bound $M=1/e$, we obtain the quantitative estimate
$\sup_{|x|\leq R}\bigl|f_{\kappa}(x)-(x)_{+}\bigr|\leq\frac{1}{e\,\kappa}\leq\eta\implies\kappa=\mathcal{O}\left(\frac{1}{\eta}\right).$
So we have
$\frac{\mathrm{SiLU}\left(\kappa(-\frac{x}{\tau})\right)}{\kappa}\xrightarrow[\kappa\to+\infty,\ \tau\to 0^{+}]{}\begin{cases}0&0\leq x\\ +\infty&0>x\end{cases}.$
$1-\frac{2}{\kappa}\text{SiLU}\left(\frac{\kappa(a-x)}{\tau}\right)\xrightarrow[\kappa\to+\infty,\ \tau\to 0^{+}]{}\begin{cases}1&a\leq x\\ -\infty&a>x\end{cases}.$
Subsequently, relation (8) leads us to
$\phi^{+}_{a,\tau,\kappa}(x)\xrightarrow[\kappa\to+\infty,\ \tau\to 0^{+}]{}\begin{cases}1&a\leq x\\ 0&a>x\end{cases}=\mathbf{1}_{[a,+\infty)}(x).$
Consequently [^1] There exist $\tau_{0}$ and $\kappa_{0}$ two strictly positive real numbers such that for all $\tau_{0}>\tau>0$ and $\kappa>\kappa_{0}$, such that $\phi^{+}_{a,\tau,\kappa}(x)-\phi^{+}_{b,\tau,\kappa}(x)\xrightarrow[]{}\mathbf{1}_{[a,+\infty)}(x)-\mathbf{1}_{[b,+\infty)}(x)=\mathbf{1}_{[a,b]}(x).$ ,
$\phi^{+}_{a,\tau,\kappa}(x)-\phi^{+}_{b,\tau,\kappa}(x)\xrightarrow[\kappa\to+\infty,\ \tau\to 0^{+}]{}\mathbf{1}_{[a,+\infty)}(x)-\mathbf{1}_{[b,+\infty)}(x)=\mathbf{1}_{[a,b]}(x).$
The proposition is proven. ∎
Figure 9: Approximation of the function $\mathbf{1}_{[1,4]}(x)$ by the function $\phi^{+}_{1,\tau,\kappa}(x)-\phi^{+}_{4,\tau,\kappa}(x)$ (Proposition 1).
We say that a function $\psi:K\to\mathbb{R}$ is a non-identically zero step function if there exists a finite partition of $K$ into pairwise disjoint intervals $I_{1},\dots,I_{m}$ with non-empty interior, and real numbers $d_{1},\dots,d_{m}$ not all zero, such that:
$\psi(x)=\sum_{j=1}^{m}d_{j}\mathbf{1}_{I_{j}}(x).$
Proposition 2 (Approximation of Step Functions).
Let $\psi:K\to\mathbb{R}$ be a non-identically zero step function:
$\psi=\sum_{j=1}^{m}d_{j}\mathbf{1}_{I_{j}},$
where the $I_{j}$ are pairwise disjoint intervals with non-empty interior, $d_{j}\in\mathbb{R}$. Then, for every $\eta>0$, there exists $N\in\mathbb{N}^{*}$, choices of $\alpha_{i},\beta_{i},c_{i},\tau,\kappa$ such that
$\left\|\psi-\sum_{i=1}^{N}c_{i}\phi_{\alpha_{i},\beta_{i},\tau,\kappa}\right\|_{\infty}<\eta.$
Proof.
We take $N=m$, $\alpha_{i},\beta_{i}$ such that $I_{i}=[\alpha_{i},\beta_{i}]$. By Proposition 1, there exist $\tau_{0}$ and $\kappa_{0}$ two strictly positive real numbers such that for all $\tau_{0}>\tau>0$ and $\kappa>\kappa_{0}$, such that
$\|\phi_{\alpha_{i},\beta_{i},\tau,\kappa}-\mathbf{1}_{I_{i}}\|_{\infty}<\delta\quad\forall i.$
We set $c_{i}=d_{i}$. Then, for $\delta=\frac{\eta}{m\cdot\max_{j}\{|d_{j}|,|c_{j}|\}}$
$\left\|\psi-\sum_{i=1}^{m}c_{i}\phi_{\alpha_{i},\beta_{i},\tau,\kappa}\right\|_{\infty}\leq\sum_{i=1}^{m}\max_{j}\{|d_{j}|,|c_{j}|\}\,\|\mathbf{1}_{I_{i}}-\phi_{\alpha_{i},\beta_{i},\tau,\kappa}\|_{\infty}<m\cdot\max_{j}\{|d_{j}|,|c_{j}|\}\cdot\delta\leq\eta.$
Proof of theorem 6.
Let $f\in C(K)$, $\epsilon>0$. Since $K$ is compact and $f$ is continuous, $f$ is uniformly continuous on $K$.
There exists $\eta>0$ such that:
$\forall x,y\in K,\ |x-y|<\eta\Rightarrow|f(x)-f(y)|<\frac{\epsilon}{2}$
We partition the compact $K$ into $m$ intervals $I_{1},\dots,I_{m}$ of diameter less than $\eta$. For each $I_{j}$, we choose $x_{j}\in I_{j}$ and define:
$\psi(x)=f(x_{j})\quad\text{for }x\in I_{j}$
Then for all $x\in K$, if $x\in I_{j}$, we have $|x-x_{j}|<\eta$ so:
$|f(x)-\psi(x)|=|f(x)-f(x_{j})|<\frac{\epsilon}{2}$
If we set $\delta=\frac{\epsilon}{4m\cdot\max_{j}\{|d_{j}|,|c_{j}|\}}$, from the triangle inequality we deduce the approximation error of the continuous function $f$
$\|f-\sum_{i=1}^{N}c_{i}\,\phi_{\alpha_{i},\beta_{i},\kappa,\tau}\|_{\infty}\leq\|f-\psi\|_{\infty}+\|\psi-\sum_{i=1}^{N}c_{i}\,\phi_{\alpha_{i},\beta_{i},\kappa,\tau}\|_{\infty}\leq\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon$
Therefore $\mathcal{F}$ is dense in $C(K)$. ∎
Corollary 3 (Fast Approximation of Continuous Functions by SiLU Neural Networks).
For $f\in C([a,b])$ with modulus of continuity $\omega_{f}$, and for $\epsilon>0$, there exists a SiLU-activated neural network, with $2$ hidden layers and width $2N$
$f_{\epsilon}(x)=\sum_{i=1}^{N}c_{i}\phi_{\alpha_{i},\beta_{i},\tau,\kappa}(x)$
such that:
$N=\mathcal{O}\left(\frac{b-a}{\omega_{f}^{-1}(\epsilon)}\right),\quad\|f-f_{\epsilon}\|_{\infty}<\epsilon$
The total number of parameters is $13N+1$.
Theorem 7 (Polynomial Approximation of Continuous Functions by SiLU Neural Networks).
Let $f\in C([-B,B])$. For every $\varepsilon\in(0,1)$, there exists a SiLU neural network $\mathcal{N}_{M,k}$ such that:
$\sup_{|x|\leq B}|f(x)-\mathcal{N}_{M,k}(x)|\leq\varepsilon$
$\mathcal{N}_{M,k}$ is a RNN with SiLU activation like in Corollary 2
To prove this theorem, we recall the following lemma:
Lemme 1.
Let $f$ be a continuous function from $[a,b]$ to $\mathbb{R}$. For every $\delta>0$, there exists a polynomial function $P_{f}$ with real coefficients such that for all $x$ in $[a,b]$, $|f(x)-P_{f}(x)|\leq\delta$.
Proof.
Let $f$ be a $C^{0}$ function on [B,B][-B,B]. Lemma 1 guarantees the existence of a polynomial function $P_{f}$ with real coefficients such that for all $x$ in [B,B][-B,B], $|f(x)-P_{f}(x)|\leq\delta=\frac{\varepsilon}{2}\leq\frac{1}{2}$. According to Theorem 4, there exists a SiLU neural network $\mathcal{N}_{M,k}$ such that
$\sup_{|x|\leq B}|P_{f}(x)-\mathcal{N}_{M,k}(x)|\leq C_{M}\cdot\max_{m}|a_{m}|\cdot\omega_{4}^{-2k}$
$\sup_{|x|\leq B}|f(x)-\mathcal{N}_{M,k}(x)|$ $=\sup_{|x|\leq B}|f(x)-P_{f}(x)+P_{f}(x)-\mathcal{N}_{M,k}(x)|$
$\leq\delta+C_{M}\cdot\max_{m}|a_{m}|\cdot\omega_{M}^{-2k}$
If we set
$k=\left\lceil-\frac{\ln\frac{\varepsilon}{2C_{M}\cdot\max_{m}|a_{m}|}}{2\ln\omega_{M}}\right\rceil,$
we obtain
$\sup_{|x|\leq B}|f(x)-\mathcal{N}_{M,k}(x)|\leq\varepsilon.$
According to Corollary 2, $\mathcal{N}_{M,k}$ can be a RNN with SiLU activation.
The following result shows that these networks can approximate any Sobolev function with a complexity that depends on the required accuracy, and establishes the optimal convergence rates depending on the regularity of the target function and the dimension of the space. Let $B>0$, $d,n\in\mathbb{N}^{*}$, and consider the open ball of Sobolev space: Functions belonging to the Sobolev class are defined as follows.
Let $\Omega=[-B,B]^{d}$ be a bounded domain. For a multi-index $\alpha=(\alpha_{1},\dots,\alpha_{d})\in\mathbb{N}^{d}$, denote $|\alpha|=\alpha_{1}+\cdots+\alpha_{d}$ and
$D^{\alpha}f=\frac{\partial^{|\alpha|}f}{\partial x_{1}^{\alpha_{1}}\cdots\partial x_{d}^{\alpha_{d}}}.$
The Sobolev space $W^{n,\infty}(\Omega)$ is defined by
[…]
It is equipped with the norm
$\|f\|_{W^{n,\infty}(\Omega)}=\max_{|\alpha|\leq n}\|D^{\alpha}f\|_{L^{\infty}(\Omega)},$
where
$\|g\|_{L^{\infty}(\Omega)}=\operatorname{ess\,sup}_{x\in\Omega}|g(x)|.$
We consider the unit ball of this space:
$\mathcal{F}_{n,d}=\big\{f\in W^{n,\infty}([-B,B]^{d}):\|f\|_{W^{n,\infty}}\leq 1\big\}.$
Theorem 8.
For any $\varepsilon\in(0,1)$, there exists a neural network $\widetilde{f}_{\varepsilon}$ with SiLU activation function such that:
$\sup_{x\in[-B,B]^{d}}|f(x)-\widetilde{f}_{\varepsilon}(x)|\leq\varepsilon,$
with a complexity of:
$\text{size}(\widetilde{f}_{\varepsilon})=\mathcal{O}\left(\varepsilon^{-d/n}\right),\quad\text{depth}(\widetilde{f}_{\varepsilon})=\mathcal{O}\left(1\right)=c(d,n,B),$
where the implicit constants depend on $d$, $n$, and $B$.
Proof.
The construction proceeds in three main stages. First, we partition the domain $B,B]^{d}$ into $M^{d}$ small cubes of side length $h\sim\varepsilon^{1/n}$. On each cube, the target function $f$ is approximated locally by its Taylor polynomial of order n1n-1. Second, these local polynomial approximations are aggregated into a global approximant $f_{\varepsilon}$ using a smooth partition of unity $\{\psi_{\mathbf{j}}\}$ constructed from sigmoidal functions. This ensures the global error is bounded by $\|f-f_{\varepsilon}\|_{\infty}\leq\varepsilon/2$. Finally, $f_{\varepsilon}$ is implemented by a SiLU neural network $\widetilde{f}_{\varepsilon}$. Key steps include approximating the partition of unity functions, the monomials $(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\alpha}$, and their products via efficient sub-networks. A detailed error analysis shows that the total approximation error can be made less than $\varepsilon$, with the network size and depth satisfying the stated bounds. The complete proof is given in the Appendix A.6. ∎
4 Discussions and experimental results
4.1 Parameterization of the Construction
Our hierarchical construction relies on three key parameters whose optimization is crucial for performance: the scale parameter $\beta$, the recursion parameter $k$, and the symmetrization shift $a$.
Figure 10: Error evolution as a function of $\beta$ and $k$.
Figure 11: Error evolution as a function of $\beta$ and $a$.
Figure 12: Error evolution as a function of the correlation between $k$ and $a$.
Scale Parameter $\beta$
The parameter $\beta$ determines the compression rate in the recursive construction $Q_{k}(x)=\beta^{2k}f_{a}(\beta^{-k}x)$. Its optimal value results from a trade-off between convergence speed and numerical stability. A systematic analysis reveals that values that are too small (e.g., $\beta<0.01$, Fig. 10) theoretically accelerate convergence but amplify numerical instabilities, while larger values ($0.1\leq\beta\leq 0.5$) improve stability and convergence speed. The value $\beta=\mathbf{0.027}$ offers a good compromise, allowing for low values of $k$ to achieve maximum reconstruction errors on the order of $10^{-7}$.
Recursion Parameter $k$
The parameter $k$, less influenced by $a$, controls the accuracy of the approximation according to the relation $Q_{k}(x)=\beta^{2k}f_{a}(\beta^{-k}x)$. Experimentally, we observe that $k=3$ is sufficient to achieve an accuracy of $10^{-6}$ in the approximation of the function $x^{2}$ (Fig. 12), while $k=5$ allows reaching $10^{-15}$ (Fig. 10). Beyond $k=10$, the benefits in accuracy are counterbalanced by the accumulation of numerical errors, illustrating the principle of maximum effective precision in our construction.
Symmetrization Shift $a$
The parameter $a$ in the symmetrized function $g_{a}(x)=(x+a)\sigma(x+a)+(-x+a)\sigma(-x+a)-2a\sigma(a)$ influences the dominant coefficient $K(a)=2\sigma^{\prime}(a)+a\sigma^{\prime\prime}(a)$. Theoretical analysis shows that $a=0$ maximizes $K(a)$ while maintaining a well-conditioned Taylor expansion. Values of $a$ close to $\approx\pm 24.5$ reduce the maximum reconstruction error of the function $x^{2}$ (Fig. 11).
4.2 Approximation of Continuous Functions
Our experiments (Fig. 13 - 17) implemented networks of by theorems 6 and 7 to approximate the following functions:
$T_{1}(x)$ $=\log(7+x)\cos(x^{3}),$
$T_{2}(x)$ $=\frac{1}{1+e^{-x}},$
$T_{3}(x)$ $=(x,\sin(x),\cos(x)),$
Figure 13: Approximation of the function $\log(7+x)\cos(x^{3})$ by theorem 7.
Figure 14: Approximation of the function $\log(7+x)\cos(x^{3})$ by theorem 6.
Figure 15: Approximation of the sigmoid function $\frac{1}{1+e^{-x}}$ by theorem 7.
Figure 16: Approximation of the sigmoid function $\frac{1}{1+e^{-x}}$ by theorem 6.
Figure 17: Approximation of the function $x\mapsto(x,\sin(x),\cos(x))$ by theorem 7.
The proposed architectures with SiLU activation demonstrate remarkable approximation capabilities for extended classes of continuous and analytic functions. Our theoretical analysis and experimental validation establish that these models can be designed as generators of adaptive functional bases, systematically producing monomials, cross-interactions, and other fundamental functional elements through deep sequential compositions.
Figure 18: Using the architecture specified in Theorem 1 to approximate the function $x^{2}$, a comparison between SiLU and ReLU activations reveals a clear performance superiority of the SiLU network.
Based on some experimental results (figs. 14 and 16), we observe good accuracy in the approximation by neural networks with SiLU activation functions via the step functions.
Comparative Performance Analysis: SiLU vs. ReLU
Experimental implementation (figs. 18 and 19) of the architectures in theorems 1 and 3 for approximating $x^{2}$ and $xy$ highlights the crucial role of parameter initialization. SiLU networks initialized near the analytical solution converge via backpropagation to values close to the theoretical optimum, whereas ReLU networks often become trapped in suboptimal minima, yielding less accurate approximations. For Sobolev functions on $B,B]^{d}$, $O(\varepsilon^{-d/n})$ multiplication blocks are computed in parallel, with the final summation performed in a single layer, producing networks of size $O(\varepsilon^{-d/n})$ and constant depth. By contrast, Yarotsky’s ReLU construction achieves the same approximation order, but the size and depth are increased by the complexity of multiplication sub-networks, introducing an additional $\ln(1/\varepsilon)$ factor.
Figure 19: The same observations hold for the network approximating the product $f(x)\cdot g(x)$($f(x)=\cos x$ and $g(x)=\sin x\log(x^{2})$) using the architecture from Theorem 2.
5 Conclusion
This work establishes the first comprehensive theoretical framework for the approximation power of SiLU networks, demonstrating their exponential convergence for basic operations and optimal Sobolev-class rates with improved depth efficiency compared to ReLU constructions. Our constructive proofs reveal an inherent parameter-shared RNN structure for polynomial approximation while providing explicit complexity bounds. These results position SiLU as a theoretically efficient activation for smooth function representation and open key research directions: deriving sharper generalization bounds, developing structured initialization schemes for scientific machine learning, and extending the framework to broader function classes and smooth activations.
Acknowledgments
We thanks Serge Iovleff for multiple helpful comments on this preliminary version of the paper. This work was supported by the Agence Française de Développement (AFD) through the Agence Nationale de la Recherche (ANR-21-PEA2-0007), under the Partenariats with "l’Enseignement supérieur Africain" (PEA) program.
Appendix A Proof of theorems
A.1 Proof of Theorem 1
Theorem 1.
For any $B>0$ and target accuracy $\epsilon\in(0,1)$ , there exist constants $C_{1}>0$, $\omega_{1}>1$ and a family of feedforward SiLU neural networks $\{Q_{k}\}_{k\in\mathbb{N}}$, with at most one hidden layer with $3$ neurons each, such that for $k=\max\left\{\left\lceil\frac{\ln C_{1}-\ln\epsilon}{2\ln\omega_{1}}\right\rceil,\left\lceil\frac{\ln 2B}{2\ln\omega_{1}}\right\rceil\right\}$
$\sup_{|x|\leq B}|Q_{k}(x)-x^{2}|\leq\epsilon,$
where $\lceil\cdot\rceil$ is the ceiling function. Moreover, the network $Q_{k}$ has depth $\mathcal{O}(1)$ and size $\mathcal{O}(1)$.
Proof.
The SiLU function is defined by
$\text{SiLU}(x)=x\sigma(x)=\frac{x}{1+e^{-x}},\qquad\forall x\in\mathbb{R}$
and $\sigma$’s successive derivatives are given by
$\sigma^{\prime}(x)$ $=\sigma(x)(1-\sigma(x))$
$\sigma^{\prime\prime}(x)$ $=\sigma^{\prime}(x)(1-2\sigma(x))$
Since $\sigma$ is $C^{\infty}$ and, more precisely, entire (analytic on $\mathbb{C}$), it admits a Taylor expansion around any point $a\in\mathbb{R}$.
The Taylor series expansion of $\sigma(z)$ around $z=a$ is:
$\sigma(z)=\sigma(a)+\sigma^{\prime}(a)(z-a)+\frac{1}{2}\sigma^{\prime\prime}(a)(z-a)^{2}+\frac{1}{6}\sigma^{\prime\prime\prime}(a)(z-a)^{3}+\frac{1}{4!}\sigma^{(4)}(a)(z-a)^{4}+\cdots+\frac{1}{n!}\sigma^{(n)}(a)(z-a)^{n}+\mathcal{O}((z-a)^{n+1})$
To simplify notation, let $s=\sigma(a),s^{\prime}=\sigma^{\prime}(a),s^{\prime\prime}=\sigma^{\prime\prime}(a),\dots$ The Taylor series expansions of $\sigma(x+a)$ and $\sigma(-x+a)$ are
$\sigma(x+a)$ $=s+s^{\prime}x+\frac{s^{\prime\prime}}{2}x^{2}+\frac{s^{\prime\prime\prime}}{6}x^{3}+\frac{s^{(4)}}{24}x^{4}+\cdots+\frac{s^{(n)}}{n!}x^{n}+\mathcal{O}(x^{n+1})$
$\sigma(-x+a)$ $=s-s^{\prime}x+\frac{s^{\prime\prime}}{2}x^{2}-\frac{s^{\prime\prime\prime}}{6}x^{3}+\frac{s^{(4)}}{24}x^{4}+\cdots+(-1)^{n}\frac{s^{(n)}}{n!}x^{n}+\mathcal{O}(x^{n+1})$
Thus, $(x+a)\sigma(x+a)$ is computed as follows
$(x+a)\sigma(x+a)$ $=(x+a)\Big[s+s^{\prime}x+\frac{s^{\prime\prime}}{2}x^{2}+\frac{s^{\prime\prime\prime}}{6}x^{3}+\frac{s^{(4)}}{24}x^{4}+\cdots+\frac{s^{(n)}}{n!}x^{n}+\mathcal{O}(x^{n+1})\Big]$
$=as+(s+as^{\prime})x+(s^{\prime}+a\frac{s^{\prime\prime}}{2})x^{2}+(\frac{s^{\prime\prime}}{2}+a\frac{s^{\prime\prime\prime}}{6})x^{3}+(\frac{s^{\prime\prime\prime}}{6}+a\frac{s^{(4)}}{24})x^{4}$
$\qquad+\cdots+\left(\frac{s^{(n-1)}}{(n-1)!}+a\frac{s^{(n)}}{n!}\right)x^{n}+\frac{s^{n}}{n!}x^{n+1}+\mathcal{O}(x^{n+2})$
Similarly,
$(-x+a)\sigma(-x+a)$ $=as-(s+as^{\prime})x+(s^{\prime}+a\frac{s^{\prime\prime}}{2})x^{2}-(\frac{s^{\prime\prime}}{2}+a\frac{s^{\prime\prime\prime}}{6})x^{3}+(\frac{s^{\prime\prime\prime}}{6}+a\frac{s^{(4)}}{24})x^{4}$
$\qquad+\cdots+(-1)^{n}\left(\frac{s^{(n-1)}}{(n-1)!}+a\frac{s^{(n)}}{n!}\right)x^{n}+(-1)^{n+1}\frac{s^{n}}{n!}x^{n+1}+\mathcal{O}(x^{n+2})$
Let
$g_{a}(x)=(x+a)\sigma(x+a)+(-x+a)\sigma(-x+a)-2a\sigma(a),$
it follows that
$g_{a}(x)$ $=\cancel{as}+\cancel{(s+as^{\prime})x}+(s^{\prime}+a\frac{s^{\prime\prime}}{2})x^{2}\cancel{+(\frac{s^{\prime\prime}}{2}+a\frac{s^{\prime\prime\prime}}{6})x^{3}}+(\frac{s^{\prime\prime\prime}}{6}+a\frac{s^{(4)}}{24})x^{4}$
$\qquad+\cdots+\left(\frac{s^{(n-1)}}{(n-1)!}+a\frac{s^{(n)}}{n!}\right)x^{n}+\frac{s^{n}}{n!}x^{n+1}+\mathcal{O}(x^{n+2})$
$\quad+\cancel{as}\cancel{-(s+as^{\prime})x}+(s^{\prime}+a\frac{s^{\prime\prime}}{2})x^{2}\cancel{-(\frac{s^{\prime\prime}}{2}+a\frac{s^{\prime\prime\prime}}{6})x^{3}}+(\frac{s^{\prime\prime\prime}}{6}+a\frac{s^{(4)}}{24})x^{4}$
$\qquad+\cdots+(-1)^{n}\left(\frac{s^{(n-1)}}{(n-1)!}+a\frac{s^{(n)}}{n!}\right)x^{n}+(-1)^{n+1}\frac{s^{n}}{n!}x^{n+1}+\mathcal{O}(x^{n+2})\cancel{-2as}$
$=(2s^{\prime}+as^{\prime\prime})x^{2}+(\frac{s^{\prime\prime\prime}}{3}+a\frac{s^{(4)}}{12})x^{4}+\cdots+(1+(-1)^{n})\left(\frac{s^{(n-1)}}{(n-1)!}+a\frac{s^{(n)}}{n!}\right)x^{n}$
$\qquad+(1+(-1)^{n+1})\frac{s^{n}}{n!}x^{n+1}+\mathcal{O}(x^{n+2})$
$=K(a)x^{2}+r(a)x^{4}+\cdots+(1+(-1)^{n})\left(\frac{s^{(n-1)}}{(n-1)!}+a\frac{s^{(n)}}{n!}\right)x^{n}+(1+(-1)^{n+1})\frac{s^{n}}{n!}x^{n+1}+\mathcal{O}(x^{n+2})$
where
$K(a)$ $=2s^{\prime}+as^{\prime\prime}$
$=s(1-s)[2+a(1-2s)]$
and
$r(a)$ $=2\frac{s^{\prime\prime\prime}}{3}+a\frac{s^{(4)}}{12}$
$=\frac{2}{3}s(1-s)(6s^{2}-6s+1)+\frac{a}{12}s(1-s)(1-2s)(12s^{2}-12s+1)$
$=s(1-s)\Big[4s^{2}-4s+\frac{2}{3}+\frac{a}{12}(1-2s)(12s^{2}-12s+1)\Big]$
To normalize the coefficient of the monomial $x^{2}$, let
$f_{a}(x)=x^{2}+c_{4,0}(a)x^{4}+\cdots+c_{n,0}(a)x^{n}+\frac{1}{2}c_{n+1,0}(0)x^{n+1}+\mathcal{O}(x^{n+2}),$
with
$c_{n,0}(a)=\frac{(1+(-1)^{n})\left(\frac{s^{(n-1)}}{(n-1)!}+a\frac{s^{(n)}}{n!}\right)}{K(a)}=c_{2j,0}(a)=\frac{2\left(\frac{s^{(2j-1)}}{(2j-1)!}+a\frac{s^{(2j)}}{(2j)!}\right)}{K(a)},\qquad\forall n=2j\geq 4$
such that $K(a)\neq 0.$
The approximation error of $f_{a}$ is of order $x^{n+2}$. To reduce this error, we exploit the homogeneity property of the dominant term $x^{2}$ and the error term $x^{n+2}$ by composing with a scaling transformation. We recursively define the sequence of functions $\{Q_{k}\}_{k}$ for $k\geq 1$ by:
$Q_{k}(x)=\beta^{-2}\cdot Q_{k-1}(\beta x),\quad\beta<1,$
with $Q_{0}(x)=f_{a}(x)$. It follows that the function $Q_{k}$ is a sequence of $k$ compositions of functions
$Q_{k}(x)=\beta^{-2k}f_{a}(\beta^{k}x)$ (9)
$Q_{k}(x)$ $=\beta^{-2}\cdot Q_{k-1}(\beta x)$
$=\beta^{-2k}(\beta^{k}x)^{2}+\beta^{-2k}\sum_{j=2}^{\infty}c_{2j,k}(a)(\beta^{k}x)^{2j}$
$=x^{2}+\sum_{j=2}^{\infty}c_{2j,0}(a)\beta^{k(2j-2)}x^{j}$
$|Q_{k}(x)-x^{2}|\leq\sum_{j=2}^{\infty}|c_{2j,0}(a)|\beta^{k(2j-2)}|x|^{2j}$
Let $B\in\mathbb{R}^{*}_{+}$ be fixed. For $x\in[-B,B]$ and $j\geq 1$, we have
$|x|\leq B$ $\implies|x|^{j}\leq B_{j}=\max\{B,B^{j}\}$
Then
$|Q_{k}(x)-x^{2}|$ $\leq\sum_{j=2}^{\infty}|c_{2j,0}(a)|\beta^{k(2j-2)}|x|^{2j}$
$\leq\sum_{j=2}^{\infty}|c_{2j,0}(a)|\beta^{2k(j-1)}B_{2j}$
$\leq\beta^{2k}\sum_{j=2}^{\infty}|c_{2j,0}(a)|\beta^{2k(j-2)}B_{2j},\qquad\text{since }0\leq j-2\text{ and }\beta<1$
$\leq\beta^{2k}\sum_{j=2}^{\infty}|c_{2j,0}(a)|\beta^{2k(j-2)}B_{2j}$
[…]
$\leq\frac{2C^{\prime}_{a}}{K(a)}\max\left\{\frac{\beta^{2k}}{1-\beta^{2k}},\frac{\beta^{2k}}{1-\beta^{2k}B^{2}}\right\}$
$\leq\frac{4C^{\prime}_{a}}{K(a)}\beta^{2k},\qquad\text{if}\ \max\left\{\beta^{2k},\beta^{2k}B^{2}\right\}\leq 1/2$
Consequently, the sequence $\{Q_{k}\}_{k}$ converges uniformly to the function $x^{2}$ on any compact set [B,B],B>0[-B,B],B>0, If $k\geq\frac{-\ln 2-\ln B}{2\ln\beta}$. By setting $\omega_{1}^{-1}=\beta$ and
$C_{1}=\frac{4C^{\prime}_{a}}{K(a)},$
The function $Q_{k}$ approximate the function $x^{2}$ with error $\epsilon=C_{1}\beta^{2k}$,where
$k=\left\lceil\frac{\ln C_{1}-\ln\epsilon}{2\ln\omega_{1}}\right\rceil$
The error decreases exponentially with the number $k$.
$k\geq\frac{\ln C_{1}-\ln\epsilon}{2\ln\omega_{1}}=\mathcal{O}\left(\ln\frac{1}{\epsilon}\right)$
Finally,
$k\geq\max\left\{\frac{\ln C_{1}-\ln\epsilon}{2\ln\omega_{1}},\frac{\ln 2B}{2\ln\omega_{1}}\right\}$
This property is highly desirable for the efficient approximation of regular functions. The function $g_{a}$ can be implemented by a SiLU network (since $g_{a}(x)=\text{SiLU}(x+a)+\text{SiLU}(-x+a)-2\text{SiLU}(a)$). The construction of the function $Q_{k}$ requires one linear operation and one evaluation of the function $f_{a}$. Therefore, we can use a SiLU neural network with depth, number of weights, and computational units all of order $\mathcal{O}(1)$ to implement the function $Q_{k}$. Indeed,
$Q_{k}(x)=\gamma\cdot\text{SiLU}(Wx+b)$ (10)
with parameters such that the weights $W=[\beta^{k},-\beta^{k},0]^{\top}$, $\gamma=\frac{\beta^{-2k}}{K(a)}[1,1,-2]$, and the bias $b=[a,a,a]^{\top}$. ∎
A.2 Proof of Theorem 2
Theorem 2.
For every $B>0$ and target accuracy $\epsilon\in(0,1)$, there exist constants $C_{2}>0$, $\omega_{2}>1$ and a family of feedforward SiLU neural networks $\{M_{k}\}_{k\in\mathbb{N}}$ suc, with at most one hidden layer with $4$ neurons each, such that for $k\geq\max\left\{\left\lceil\frac{\ln C_{2}-\ln\epsilon}{2\ln\omega_{2}}\right\rceil,\left\lceil\frac{\ln 2B}{2\ln\omega_{2}}\right\rceil\right\}$
$\sup_{|x|,|y|\leq B}|M_{k}(x,y)-xy|\leq\epsilon$
The network $M_{k}$ has depth $O(1)$ and size $O(1)$.
Proof.
We will prove this theorem using the well-known identity for the product:
$xy=\frac{1}{4}[(x+y)^{2}-(x-y)^{2}]$
This identity allows us to express the product $xy$ as a linear combination of two squares. Thanks to Theorem 1, we know how to approximate the function $x^{2}$ using SiLU neural networks with an error that decreases exponentially with depth. Let us define the auxiliary functions $A(x,y)=x+y$ and S(x,y)=xyS(x,y)=x-y. These functions are linear and can be computed exactly by a neural network (for example, a linear layer without activation).
Let $\{Q_{k}\}_{k\in\mathbb{N}^{*}}$ be the family of SiLU neural networks given by Theorem 1 that approximates the function $x^{2}$ with a uniform error on [2B,2B][-2B,2B] bounded by $C_{1}\omega^{-2k}$, with $\omega>1$ and $C_{1}>0$. Now define the network $M_{k}$ by
$M_{k}(x,y)=\frac{1}{4}[Q_{k}(A(x,y))-Q_{k}(S(x,y))],$
with the network $Q_{k}$ defined in (9).
This network is a composition of linear networks (to compute $A$ and $S$) and the networks $Q_{k}$ (to approximate the squares). The final output is a linear combination of the results. Let us estimate the approximation error of the function $xy$ by $M_{k}(x,y)$.
$|M_{k}(x,y)-xy|$ $=\left|\frac{1}{4}\left[Q_{k}(x+y)-Q_{k}(x-y)\right]-\frac{1}{4}\left[(x+y)^{2}-(x-y)^{2}\right]\right|$
$\leq\frac{1}{4}\left|Q_{k}(x+y)-(x+y)^{2}\right|+\frac{1}{4}\left|Q_{k}(x-y)-(x-y)^{2}\right|.$
Since $|x+y|\leq 2B$ and $|x-y|\leq 2B$, we can apply Theorem 1. There exist $C_{1}>0,C^{\prime}_{1}>0,\omega_{1}>1$ and $\omega^{\prime}_{1}>1$ such that
$|Q_{k}(x+y)-(x+y)^{2}|\leq C_{1}\omega_{1}^{-2k},\qquad$ $|Q_{k}(x-y)-(x-y)^{2}|\leq C^{\prime}_{1}{\omega^{\prime}}_{1}^{-2k}$
Thus,
$|M_{k}(x,y)-xy|$ $\leq C_{2}\omega_{2}^{-2k},\qquad\text{for all }|x|,|y|\leq B$
Therefore,
$\sup_{|x|,|y|\leq B}|M_{k}(x,y)-xy|$ $\leq C_{2}\omega_{2}^{-2k},$
with $C_{2}=\frac{C_{1}+C^{\prime}_{1}}{4}$ and $\omega_{2}=2\min\{\omega_{1},\omega^{\prime}_{1}\}$.The NN $M_{k}(x,y)$ approximate the function $(x,y)\mapsto x\cdot y$ with error $\epsilon=C_{2}w_{2}^{-2k}$, where
$k=\left\lceil\frac{\ln C_{2}-\ln\epsilon}{2\ln w_{2}}\right\rceil.$
Finally $k\geq\max\left\{\left\lceil\frac{\ln C_{2}-\ln\epsilon}{2\ln\omega_{2}}\right\rceil,\left\lceil\frac{\ln 2B}{2\ln\omega_{2}}\right\rceil\right\}$.
Let us now analyze the complexity of the network. The network $M_{k}$ is constructed as follows:
  • The computation of $A(x,y)=x+y$ and S(x,y)=xyS(x,y)=x-y each requires 2 linear layers with 2 inputs and 1 output (without an activation function),
  • The implementation of $Q_{k}(A(x,y))$ and $Q_{k}(S(x,y))$ according to Theorem 1 requires two SiLU networks, each of depth and size $\mathcal{O}(1)$,
  • $M_{k}$ is obtained via a linear layer with 2 inputs (without activation) and 1 output.
Since the two applications of $Q_{k}$ are independent, they can be computed in parallel. Thus, the total depth of $M_{k}$ is $\mathcal{O}(1)$ plus a constant (for the initial and final linear layers). The total size is $\mathcal{O}(1)$. ∎
Therefore, one can use a SiLU neural network with a depth, number of weights, and computational units all of the order of $\mathcal{O}(1)$ to implement the function $M_{k}$. Indeed,
$M_{k}(x,y)=\gamma\cdot SiLU(Wz+b)$ (11)
with parameters such as the weights and bias
$W=\beta^{k}\begin{bmatrix}+1&+1\\ -1&-1\\ +1&-1\\ -1&+1\end{bmatrix},\quad\gamma=\frac{\beta^{-2k}}{4K(a)}\begin{bmatrix}+1\\ +1\\ -1\\ -1\end{bmatrix}^{\top},\quad b=[a,a,a,a]^{\top},\quad z=[x,y]^{\top}$
A.3 Proof of Theorem 3
Theorem 3.
For every $B>0,m\geq 3,$ and target accuracy $\epsilon\in(0,1)$, there exist constants $C_{m}>0$, $\omega_{m}>1$ and a family of feedforward SiLU neural networks $\{P_{m,k}\}_{k\in\mathbb{N}}$, with at most $2m$ hidden layers with $4$ neurons each, such that for $k\geq\max\left\{\left\lceil\frac{\ln C_{m}-\ln\epsilon}{2\ln\omega_{m}}\right\rceil,\left\lceil\frac{\ln 2B}{2\ln\omega_{m}}\right\rceil\right\}$
$\sup_{|x|\leq B}|P_{m,k}(x)-x^{m}|\leq\epsilon$
The network $P_{m,k}$ has depth $O(m)$ and size $O(1)$.
Proof.
The proof proceeds by induction on the degree $m$.
  • For $m=0$: The constant function $x^{0}=1$ can be represented exactly by a neural network with no parameters (constant output 1). Thus, we set $P_{0,k}(x)=1$ for all $k\geq 1$, resulting in zero error. The depth is $O(1)$ and the size is $O(1)$.
  • For $m=1$: The identity function $x^{1}=x$ can be represented exactly by a linear network (linear output). We set $P_{1,k}(x)=x$ for all $k$, resulting in zero error. The depth is $O(1)$ and the size is $O(1)$.
  • For $m=2$: By Theorem 1, there exist constants $C_{1}>0$, $\omega_{1}>1$, and a family of SiLU neural networks $\{Q_{k}\}_{k\in\mathbb{N}^{*}}$ such that for all $k$:
    $\sup_{|x|\leq B}|Q_{k}(x)-x^{2}|\leq C_{1}\cdot\omega_{1}^{-2k}.$
    We define $P_{2,k}(x)=Q_{k}(x)$. The depth is $\mathcal{O}(1)$ and the size is $\mathcal{O}(1)$. Note that for $m=2$, the size is $O(1)$.
Assume that for some $m\geq 2$, there exists a family of SiLU networks $\{P_{m,k}\}_{k\in\mathbb{N}}$ such that:
$|P_{m,k}(x)-x^{m}|\leq C_{m}\omega_{m}^{-2k},$
with depth and size $\mathcal{O}(m)$.
We aim to approximate $x^{m+1}=x^{m}\cdot x$. We use the multiplier network $M_{k}$ from the previous theorem 3 (which approximates the product) with inputs $P_{m,k}(x)$ (approximating $x^{m}$) and $x$ (exact). However, the multiplier $M_{k}$ is defined for bounded inputs. Assume that since $|x|\leq B$, $x^{m}$ is bounded on $D_{m,B},D_{m,B}]$. The approximation of $x^{m}$, $P_{m,k}(x)$, is also bounded on $D_{m,B},D_{m,B}]$, for all $|x|\leq B$ and all $k$ (due to uniform approximation). By taking $B^{\prime}=B+D_{m,B}$, we can apply $M_{k}$ on the domain $B^{\prime},B^{\prime}]$ to approximate $x^{m+1}$. We define:
$P_{m+1,k}(x)=M_{k}(P_{m,k}(x),x).$
Now, estimate the error:
$\sup_{|x|\leq B}\left|P_{m+1,k}(x)-x^{m+1}\right|$ $=\sup_{|x|\leq B}\left|M_{k}(P_{m,k}(x),x)-x^{m}\times x\right|$
$\leq\sup_{|x|\leq B}\left|M_{k}(P_{m,k}(x),x)-P_{m,k}(x)\times x\right|+\sup_{|x|\leq B}\left|P_{m,k}(x)\times x-x^{m}\times x\right|$
$\leq C_{2}\omega_{2}^{-2k}+BC_{m}\omega_{m}^{-2k}.$
The last inequality holds because, by the multiplier theorem, there exist $C_{2}>0$ and $\omega_{2}>1$ such that:
$\sup_{|x|\leq B}\left|M_{k}(P_{m,k}(x),x)-P_{m,k}(x)\cdot x\right|\leq C_{2}\omega_{2}^{-2k}.$
Let $\omega_{m+1}=2\min\{\omega_{2},\omega_{m}\}$ and $C_{m+1}=C_{2}+B^{\prime}C_{m}$. Then there exists $C_{m+1}>0$ such that:
$\sup_{|x|\leq B}\left|P_{m+1,k}(x)-x^{m+1}\right|\leq C_{m+1}\omega_{m+1}^{-2k}.$
The network $P_{m,k}$ has depth $\mathcal{O}(m)$ and size $\mathcal{O}(1)$. So, the NN $P_{m,k}(x)$ approximate the monomial function $x\mapsto x^{m}$ with error $\epsilon=C_{m}\omega_{m}^{-2k}$,where
$k=\left\lceil\frac{\ln C_{m}-\ln\epsilon}{2\ln\omega_{m}}\right\rceil.$
$P_{m,k}(x)=\gamma_{m}\cdot\text{SiLU}\left(W_{m}z_{m}+b_{m}\right),$
with parameters such as weights and bias:
$W_{i}=\beta^{k}\begin{bmatrix}+1&+1\\ -1&-1\\ +1&-1\\ -1&+1\end{bmatrix},\quad\gamma_{i}=\frac{\beta^{-2k}}{4K(a)}\begin{bmatrix}+1\\ +1\\ -1\\ -1\end{bmatrix}^{\top},\quad b_{i}=[a,a,a,a]^{\top},\quad z_{i}=[x,\gamma_{i-1}\cdot\text{SiLU}(Wz_{i-1}+b_{i-1})]^{\top},$
and $z_{1}=[x,1]^{\top}$. ∎
A.4 Proof of Theorem 4
Theorem 4.
Let $N(x)=\sum_{m=0}^{M}a_{m}x^{m}$ be a polynomial of degree $M$. For every $B>0$ and target accuracy $\epsilon\in(0,1)$, there exist constants $C_{M}>0$, $\omega_{M}>1$ and a family of SiLU neural network $\{\mathcal{N}_{M,k}\}_{k\in\mathbb{N}}$, with at most $2M$ hidden layers with $4$ neurons each, such that for $k=\left\{\left\lceil\frac{\ln(C_{M}\cdot\max_{m}|a_{m}|)-\ln\epsilon}{2\ln\omega_{M}}\right\rceil,\left\lceil\frac{\ln 2B}{2\ln\omega_{M}}\right\rceil\right\}$
$\sup_{|x|\leq B}|\mathcal{N}_{M,k}(x)-N(x)|\leq\epsilon$
The network has depth $O(M)$ and size $O(1)$.
Proof.
Let $B>0$. According to the previous theorem, for each $m\in\{0,1,\dots,M\}$, there exists a SiLU neural network $P_{m,k}$ such that:
$\sup_{|x|\leq B}|P_{m,k}(x)-x^{m}|\leq C_{m}\omega_{3}^{-2k},\qquad\text{for all }x\in[-B,B]$
with depth $O(m)$ and size $O(1)$.
We define the network $\mathcal{N}_{M,k}$ by the linear combination:
$\mathcal{N}_{M,k}(x)=\sum_{m=0}^{M}a_{m}P_{m,k}(x).$
This is indeed a SiLU network, as a linear combination of SiLU networks remains a SiLU network (by adding a linear output layer).
Let us bound the error:
$\sup_{|x|\leq B}|\mathcal{N}_{M,k}(x)-N(x)|=\sup_{|x|\leq B}\left|\sum_{m=0}^{M}a_{m}(P_{m,k}(x)-x^{m})\right|\leq\sum_{m=0}^{M}|a_{m}|\sup_{|x|\leq B}|P_{m,k}(x)-x^{m}|.$
Since $|a_{m}|\leq\max_{0\leq m\leq M}|a_{m}|$ for all $m$, we obtain:
$\sup_{|x|\leq B}|\mathcal{N}_{M,k}(x)-N(x)|\leq\max_{m}|a_{m}|\sum_{m=0}^{M}\sup_{|x|\leq B}|P_{m,k}(x)-x^{m}|\leq\max_{m}|a_{m}|\cdot\sum_{m=0}^{M}C_{m}\cdot\omega_{m}^{-2k}.$
Thus,
$|\mathcal{N}_{M,k}(x)-N(x)|\leq\max_{m}|a_{m}|\cdot C_{m}(M+1)\cdot\omega_{m}^{-2k}.$
By setting $C_{M}=C_{m}(M+1)$ and $\omega_{M}=\omega_{m}$, we have:
$\sup_{|x|\leq B}|\mathcal{N}_{M,k}(x)-N(x)|\leq C_{M}\omega_{M}^{-2k}\max_{m}|a_{m}|.$
Let us analyze the complexity of the network:
  • Each $P_{m,k}$ has depth $O(m)$.
  • $P_{M,k}$ is an RNN of depth $O(M)$ and width $O(1)$ that, given the same input $x$, computes in its $m$-th hidden layer an approximation $P_{m,k}$ of the monomial $x^{m}$, for all $m\leq M$.
  • The final linear combination is a layer with $M+1$ inputs and 1 output.
Thus, the total depth is dominated by that of the deepest $P_{m,k}$, i.e., $O(M)$. The total size is the sum of the sizes of the $P_{M,k}$. The SiLU neural network, with depth $\mathcal{O}(M)$, and size $\mathcal{O}(1)$, well approximates the function $N(x)$. Indeed, for all $m\geq 2$:
$\mathcal{N}_{M,k}=\mathbf{c}^{\top}\mathbf{P}_{k}(x)$
with parameters such as the weights:
$\mathbf{P}_{k}(x)=\begin{bmatrix}P_{0,k}(x)\\ P_{1,k}(x)\\ \vdots\\ P_{m,k}(x)\end{bmatrix},\quad\mathbf{c}=[c_{0},c_{1},\cdots,c_{m}]^{\top},$
where $P_{0,k}(x)=1$ and $P_{1,k}(x)=x$. It’s clear that the NN $\mathcal{N}_{M,k}$ approximate the polynomial function $\sum_{m=0}^{M}a_{m}x^{m}$ with error $\epsilon=C_{M}\cdot\max_{m}|a_{m}|\omega_{M}^{-2k}$,where
$k=\left\lceil\frac{\ln(C_{M}\cdot\max_{m}|a_{m}|)-\ln\epsilon}{2\ln\omega_{M}}\right\rceil.$
A.5 Proof of Theorem 5
Theorem 5.
For any $B>0$ and target accuracy $\epsilon\in(0,1)$, there exist constants $C_{1}>0$, $\omega_{1}>1$ and a family of feedforward SiLU neural networks $\{Q^{m}_{k}\}_{k\in\mathbb{N}}$, with one hidden layer with $m+1$ neurons each, such that for $k=\max\left\{\left\lceil\frac{\ln mB}{\ln\omega_{1}}\right\rceil,\left\lceil\frac{\ln C_{1,m}-\ln\epsilon}{\ln\omega_{1}}\right\rceil\right\}$
$\sup_{|x|\leq B}|Q^{m}_{k}(x)-x^{m}|\leq\epsilon,$
Moreover, the network $Q^{m}_{k}$ has depth $\mathcal{O}(1)$ and size $\mathcal{O}(m)$.
Proof.
We will now revisit the entire general proof for the construction of a monomial $x^{m}$ by combining SiLU functions, following the rigorous logic we used to approximate $x^{2}$, but in the general case.
Recall that:
$\mathrm{SiLU}(x)=x\sigma(x),\qquad\sigma(x)=\frac{1}{1+e^{-x}}.$
For any $a\in\mathbb{R}$ and any integer $m\geq 1$, define
$g_{a,m}(x)=\sum_{j=0}^{m}(-1)^{m-j}\binom{m}{j}\mathrm{SiLU}\left((j-\tfrac{m}{2})x+a\right).$
Rewrite
$\mathrm{SiLU}\left((j-\tfrac{m}{2})x+a\right)=\left((j-\tfrac{m}{2})x+a\right)\sigma\left((j-\tfrac{m}{2})x+a\right)$
Since the function $\sigma$ is analytic on $\mathbb{R}$, we can expand it in a Taylor series around $x=0$
$\sigma\left(x\left(j-\frac{m}{2}\right)+a\right)=\sum_{n=0}^{\infty}\frac{\left(j-\frac{m}{2}\right)^{n}}{n!}\sigma^{(n)}(a)x^{n}$
Multiplying by $\left(x\left(j-\frac{m}{2}\right)+a\right)$, we obtain
$\displaystyle\left(x\left(j-\frac{m}{2}\right)+a\right)\sigma\left(x\left(j-\frac{m}{2}\right)+a\right)$ $=\left(x\left(j-\frac{m}{2}\right)+a\right)\sum_{n=0}^{\infty}\frac{\left(j-\frac{m}{2}\right)^{n}}{n!}\sigma^{(n)}(a)x^{n}$
$=_{u=j-\frac{m}{2}}[xu+a]\cdot\sum_{n=0}^{\infty}\frac{u^{n}}{n!}\sigma^{(n)}(a)x^{n}$
$=xu\cdot\sum_{n=0}^{\infty}\frac{u^{n}}{n!}\sigma^{(n)}(a)x^{n}+a\cdot\sum_{n=0}^{\infty}\frac{u^{n}}{n!}\sigma^{(n)}(a)x^{n}$
$\displaystyle\left(x\left(j-\frac{m}{2}\right)+a\right)\sigma\left(x\left(j-\frac{m}{2}\right)+a\right)$ $=\sum_{n=0}^{\infty}\frac{u^{n+1}}{n!}\sigma^{(n)}(a)x^{n+1}+\sum_{n=0}^{\infty}\frac{au^{n}}{n!}\sigma^{(n)}(a)x^{n}$
$=\sum_{m=1}^{\infty}\frac{u^{m}}{(m-1)!}\sigma^{(m-1)}(a)x^{m}+\sum_{n=0}^{\infty}\frac{au^{n}}{n!}\sigma^{(n)}(a)x^{n}$
$=\sum_{n=0}^{\infty}\frac{au^{n}}{n!}\sigma^{(n)}(a)x^{n}+\sum_{n=1}^{\infty}\frac{u^{n}}{(n-1)!}\sigma^{(n-1)}(a)x^{n}$
$=a\sigma(a)+\sum_{n=1}^{\infty}\left[\frac{au^{n}}{n!}\sigma^{(n)}(a)+\frac{u^{n}}{(n-1)!}\sigma^{(n-1)}(a)\right]x^{n}$
$=a\sigma(a)+\sum_{n=1}^{\infty}\frac{u^{n}}{n!}\left[a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)\right]x^{n}$
$=a\sigma(a)+\sum_{n=1}^{\infty}\frac{\left(j-\frac{m}{2}\right)^{n}}{n!}\left[a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)\right]x^{n}$
Substituting into the expression
$g_{a,m}(x)=\sum_{j=0}^{m}(-1)^{m-j}\binom{m}{j}\mathrm{SiLU}\left((j-\tfrac{m}{2})x+a\right),$
it follows that
$g_{a,m}(x)$ $=\sum_{j=0}^{m}(-1)^{m-j}\binom{m}{j}\mathrm{SiLU}\left((j-\tfrac{m}{2})x+a\right)$
$=a\sigma(a)\underbrace{\sum_{j=0}^{m}(-1)^{m-j}\binom{m}{j}}_{=(1-1)^{m}=0}+\sum_{j=0}^{m}(-1)^{m-j}\binom{m}{j}\sum_{n=1}^{\infty}\frac{\left(j-\frac{m}{2}\right)^{n}}{n!}\left[a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)\right]x^{n}$
$=\sum_{n=1}^{\infty}\sum_{j=0}^{m}(-1)^{m-j}\binom{m}{j}\frac{\left(j-\frac{m}{2}\right)^{n}}{n!}\left[a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)\right]x^{n}$
$=\sum_{n=1}^{\infty}\sum_{j=0}^{m}(-1)^{m-j}\binom{m}{j}\frac{\left(j-\frac{m}{2}\right)^{n}}{n!}\left[a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)\right]x^{n}$
Acoording to Appendix B, if $k=m,$
$\Delta^{m}x^{\underline{m}}=m!\,x^{\underline{0}}=m!$
and if $k<m,$ then
$\Delta^{m}x^{k}=\Delta^{m-k}\Delta^{k}x^{k}=\Delta^{m-k}k!=0$
$g_{a,m}(x)$ $=\sum_{n=1}^{\infty}\frac{\Delta^{m}p_{n}(0)}{n!}\left[a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)\right]x^{n},\qquad p_{n}(x)=\left(x-\frac{m}{2}\right)^{n}$
$=\frac{\overbrace{\Delta^{m}p_{m}(0)}^{=m!}}{m!}\left[a\sigma^{(m)}(a)+m\sigma^{(m-1)}(a)\right]x^{m}+\sum_{n=m+1}^{\infty}\frac{\Delta^{m}p_{n}(0)}{n!}\left[a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)\right]x^{n}$
$=\left[\underbrace{a\sigma^{(m)}(a)+m\sigma^{(m-1)}(a)}_{=K_{m}(a)}\right]x^{m}+\sum_{n=m+1}^{\infty}\frac{\Delta^{m}p_{n}(0)}{n!}\left[a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)\right]x^{n}$
$=K_{m}(a)x^{m}+\sum_{n=m+1}^{\infty}\frac{\Delta^{m}p_{n}(0)}{n!}\left[a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)\right]x^{n},\qquad K_{m}(a)\neq 0$
Subsequently, we define
$f_{a,m}(x)=\frac{1}{K_{m}(a)}g_{a,m}(x)=x^{m}+\sum_{n=m+1}^{\infty}\frac{\Delta^{m}p_{n}(0)}{n!K_{m}(a)}\left[a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)\right]x^{n}$
such that
$K_{m}(a)\neq 0.$
This is thus an approximation of $x^{m}$ of order $m+1$ constructed solely from combinations of SiLU. For $0<\beta<1$, define the sequence of functions $\{Q^{m}_{k}\}_{k\geq 0}$ by recurrence:
$Q^{m}_{0}(x):=f_{a,m}(x),\quad Q^{m}_{k}(x):=\beta^{-m}Q_{k-1}(\beta x).$
This is exactly analogous to the construction for $x^{2}$. The factor $\beta^{-m}$ cancels the effect of the main $x^{m}$, and the contraction $\beta^{k}$ reduces the higher powers.
We show by induction that:
$Q^{m}_{k}(x)=\beta^{-mk}f_{a,m}(\beta^{k}x)=\beta^{-mk}\left[(\beta^{k}x)^{m}+\sum_{n=m+1}^{\infty}\frac{\Delta^{m}p_{n}(0)}{n!K_{m}(a)}\left[a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)\right](\beta^{k}x)^{n}\right].$
We thus obtain:
$Q^{m}_{k}(x)=x^{m}+\sum_{n=m+1}^{\infty}\frac{\Delta^{m}p_{n}(0)}{n!K_{m}(a)}\left[a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)\right]\beta^{k(n-m)}x^{n}.$
We will calculate the maximum approximation error of the monomial $x^{m}$ by the network $Q^{m}_{k}.$ Thus, the difference $Q^{m}_{k}(x)-x^{m}$ is written
$Q^{m}_{k}(x)-x^{m}=\sum_{n=m+1}^{\infty}A_{n}\,\beta^{k(n-m)}x^{n}$
with
$A_{n}=\frac{\Delta^{m}p_{n}(0)}{n!\,K_{m}(a)}\left[a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)\right],\quad p_{n}(j)=\left(j-\frac{m}{2}\right)^{n}.$
We know that $\Delta^{m}p_{n}(0)=\sum_{j=0}^{m}(-1)^{m-j}\binom{m}{j}\left(j-\frac{m}{2}\right)^{n}$ and that $\max_{0\leq j\leq m}\left|j-\frac{m}{2}\right|=\frac{m}{2}.$ Then
$\left|j-\frac{m}{2}\right|\leq\frac{m}{2}.$
Consequently
$|\Delta^{m}p_{n}(0)|\leq\sum_{j=0}^{m}\binom{m}{j}\left(\frac{m}{2}\right)^{n}=\left(\frac{m}{2}\right)^{n}\sum_{j=0}^{m}\binom{m}{j}=\left(\frac{m}{2}\right)^{n}2^{m}.$
So
$|\Delta^{m}p_{n}(0)|\leq 2^{m}\left(\frac{m}{2}\right)^{n}.$
According to (13) of Appendix C, there exists a constant $C_{a}$ such that
$|a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)|\leq|a|C_{a}n!+nC_{a}(n-1)!=C_{a}(|a|n!+n!)=C_{a}(|a|+1)n!.$
So
$|a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)|\leq C^{\prime}_{a}n!.$
The upper bound of $|A_{n}|$ follows
$|A_{n}|\leq\frac{2^{m}(m/2)^{n}}{n!\,|K_{m}(a)|}\cdot C^{\prime}_{a}n!=\frac{2^{m}C^{\prime}_{a}}{|K_{m}(a)|}\left(\frac{m}{2}\right)^{n}.$
So
$|A_{n}|\leq C_{1}\left(\frac{m}{2}\right)^{n},\quad C_{1}=\frac{2^{m}C^{\prime}_{a}}{|K_{m}(a)|}.$
We can also bound $|Q^{m}_{k}(x)-x^{m}|$ for $|x|\leq B\leq 1$ by
$|Q^{m}_{k}(x)-x^{m}|\leq\sum_{n=m+1}^{\infty}C_{1}\left(\frac{m}{2}\right)^{n}\beta^{k(n-m)}|x|^{n}.$
For $|x|\leq B\leq\implies|x|^{n}\leq B_{n}=\max\{1,B^{n}\}$
$|Q^{m}_{k}(x)-x^{m}|\leq C_{1}\sum_{n=m+1}^{\infty}\left(\frac{m}{2}\right)^{n}\beta^{k(n-m)}B_{n}=C_{1}\beta^{-mk}\sum_{n=m+1}^{\infty}\left(\frac{mB\beta^{k}}{2}\right)^{n}.$
Let $y=\frac{mB\beta^{k}}{2}$.
For $k$ large enough such that $y<1/2$:
$\sum_{n=m+1}^{\infty}y^{n}=y^{m+1}\sum_{n=0}^{\infty}y^{n}=\frac{y^{m+1}}{1-y}.$
Therefore
$|Q^{m}_{k}(x)-x^{m}|\leq C_{1}\beta^{-mk}\cdot\frac{y^{m+1}}{1-y}.$
But $\beta^{-mk}y^{m+1}=\beta^{-mk}\left(\frac{mB\beta^{k}}{2}\right)^{m+1}=\beta^{-mk}\beta^{k(m+1)}\left(\frac{mB}{2}\right)^{m+1}=\beta^{k}\left(\frac{mB}{2}\right)^{m+1}$.
Thus:
$|Q^{m}_{k}(x)-x^{m}|\leq C_{1}\left(\frac{mB}{2}\right)^{m+1}\frac{\beta^{k}}{1-\frac{mB\beta^{k}}{2}}.$
For $k\geq k_{0}$ such that $k_{0}=\frac{-\ln mB}{\ln\beta}$, we have $y<1/2$ and thus
$\sup_{x\in[-B,B]}|Q^{m}_{k}-x^{m}|\leq 2C_{1}\left(\frac{mB}{2}\right)^{m+1}\beta^{k}.$
So
$\sup_{x\in[-B,B]}|Q^{m}_{k}-x^{m}|\leq C_{1,m}\omega_{1}^{-k}$
with $C_{1,m}=2C_{1}\left(\frac{mB}{2}\right)^{m+1}=\frac{C^{\prime}_{a}(mB)^{m+1}}{|K_{m}(a)|}$ and $\beta=\omega_{1}^{-1}.$
The convergence is thus uniform on every compact set, exponentially fast in $k$. For the maximum approximation error to be $\epsilon,$ we need $k\geq-\frac{\ln\epsilon-\ln C_{1,m}}{\ln\omega_{1}}=k_{1}.$ Adding the condition $y<1$ for the convergence of the series, we take $k=\max\{k_{0},k_{1}\}.$ So, we can use a SiLU neural network with one hidden layer, and $m+1$ neurons per layer to implement the NN $Q^{m}_{k}$. Indeed,
$Q^{m}_{k}(x)=\gamma_{m}\cdot\text{SiLU}(W_{m}x+b_{m})$ (12)
with parameters such that the weights
$W=\beta^{k}\,[w_{0},w_{1},\dots,w_{m}]^{\top},\gamma=\frac{\beta^{-mk}}{K_{m}(a)}[c_{0},c_{1},\dots,c_{m}],$
and the bias
$b=a\mathbf{1}_{m+1}.$
with
$w_{j}=j-\frac{m}{2},\quad c_{j}=(-1)^{m-j}\binom{m}{j}$,
and where $\mathbf{1}_{m+1}$ is the column vector of size $m+1$ containing ones. ∎
A.6 Proof of Theorem 8
Theorem 8.
For any $\varepsilon\in(0,1)$, there exists a neural network $\widetilde{f}_{\varepsilon}$ with SiLU activation function such that:
$\sup_{x\in[-B,B]^{d}}|f(x)-\widetilde{f}_{\varepsilon}(x)|\leq\varepsilon,$
with a complexity of:
$\text{size}(\widetilde{f}_{\varepsilon})=O\left(\varepsilon^{-d/n}\right),\quad\text{depth}(\widetilde{f}_{\varepsilon})=O\left(1\right)=c(d,n,B),$
where the implicit constants depend on $d$, $n$, and $B$.
Proof.
The proof proceeds in three key steps: a discretization of the space into small cubes where the function is locally approximated by its Taylor polynomial; an aggregation of these local approximations into a global and regular function via sigmoidal bump functions; and finally, the efficient implementation of this construction using a SiLU neural network, whose complexity we analyze.
Partition the domain $B,B]^{d}$ into $M^{d}$ cubes of side length $h=\frac{2B}{M}$. Each cube $C_{\mathbf{j}}$ ($j=1,\ldots,M^{d}$) is of the form:
$C_{\mathbf{j}}=\prod_{i=1}^{d}\left[c_{j_{i}}-\frac{h}{2},c_{j_{i}}+\frac{h}{2}\right]=\prod_{i=1}^{d}\left[\alpha_{j_{i}},\beta_{j_{i}}\right],\qquad\mathbf{j}=(j_{1},j_{2},\cdots,j_{d})\in\{1,2,\cdots,M\}^{d}=M_{d}$
where $\mathbf{c}_{\mathbf{j}}=(c_{j_{1}},c_{j_{2}},\cdots,c_{j_{d}})$ is the center of $C_{\mathbf{j}}$. For each cube $C_{\mathbf{j}}$, define a function $\psi_{\mathbf{j}}:\mathbb{R}^{d}\to[0,1]$ that is approximately 1 on $C_{\mathbf{j}}$ and decays exponentially outside. Define
$\psi_{\mathbf{j}}(\mathbf{x})=\prod_{i=1}^{d}\mathbf{1}_{[\alpha_{j_{i}},\beta_{j_{i}}]}(x_{i}),$
Figure 20: Partition of unity by the functions $\{\psi_{j}\}_{j=1}^{9}$ from Theorem 8 with $d=1$ and $M=9$. The sum is shown as a black dashed line.
It follows that
$\sum_{\mathbf{j}}\psi_{\mathbf{j}}(\mathbf{x})\equiv 1,\qquad\mathbf{x}\in[-B,B]^{d}.$
For each cube $C_{\mathbf{j}}$ and for any $f\in\mathcal{F}_{n,d}$, we consider the Taylor polynomial of order n1n-1 at the point $\mathbf{x}=\mathbf{c}_{\mathbf{j}}$:
$T_{\mathbf{j}}(\mathbf{x})=\sum_{|\bm{\alpha}|\leq n-1}\frac{D^{\bm{\alpha}}f(\mathbf{c}_{\mathbf{j}})}{\bm{\alpha}!}(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\bm{\alpha}}.$
We define the global approximant of the function $f$
$f_{\varepsilon}(\mathbf{x})=\sum_{\mathbf{j}}\psi_{\mathbf{j}}(\mathbf{x})T_{\mathbf{j}}(\mathbf{x}).$
$\left|f(\mathbf{x})-f_{\varepsilon}(\mathbf{x})\right|$ $=\left|\sum_{\mathbf{j}}\psi_{\mathbf{j}}(\mathbf{x})\left(f(\mathbf{x})-T_{\mathbf{j}}(\mathbf{x})\right)\right|$
$\leq\sum_{\mathbf{j}\in C_{\mathbf{j}}}\left|\left(f(\mathbf{x})-T_{\mathbf{j}}(\mathbf{x})\right)\right|,\qquad\text{since}\ \psi_{\mathbf{j}}(\mathbf{x})\leq 1$
$\leq 2^{d}\max_{\mathbf{j}\in C_{\mathbf{j}}}\left|\left(f(\mathbf{x})-T_{\mathbf{j}}(\mathbf{x})\right)\right|$
$\leq 2^{d}\frac{h^{n}}{n!2^{n}}\max_{|\bm{\alpha}|=n}\text{ess sup}_{\mathbf{x}\in[-B,B]^{d}}\left|D^{\bm{\alpha}}f(\mathbf{x})\right|$
$\leq 2^{d}\frac{h^{n}}{n!2^{n}}$
The last three inequalities hold because respectively: in each direction $\mathbf{j}$, there are at most $2$ cubes bounding $C_{\mathbf{j}}$; for $\mathbf{x}\in C_{\mathbf{j}}$, we have $|x_{i}-c_{j_{i}}|\leq\frac{h}{2}$ for all $i$, hence
$|\mathbf{x}-\mathbf{c}_{\mathbf{j}}|^{\bm{\alpha}}=\prod_{i=1}^{d}|x_{i}-c_{j_{i}}|^{\alpha_{i}}\leq\left(\frac{h}{2}\right)^{|\bm{\alpha}|}=\left(\frac{h}{2}\right)^{n},$
and $\text{ess sup}_{\mathbf{x}\in[-B,B]^{d}}\left|D^{\bm{\alpha}}f(\mathbf{x})\right|\leq 1.$
We want $\|f-f_{\varepsilon}\|_{\infty}\leq\frac{\varepsilon}{2}$. It suffices that
$2^{d}\frac{h^{n}}{n!2^{n}}\leq\frac{\varepsilon}{2}\quad\Rightarrow\quad h\leq 2\left(\frac{(n!)\varepsilon}{2^{d+1}}\right)^{1/n}.$
Since $h=\frac{2B}{M}$, we choose:
$M=\left\lceil\frac{2B}{h}\right\rceil=\left\lceil B\left(\frac{2^{d+1}}{(n!)\varepsilon}\right)^{1/n}\right\rceil.$
Thus, $M=O(\varepsilon^{-1/n})$ and $N_{\text{cubes}}=M^{d}=O(\varepsilon^{-d/n})$. In total, we have constructed the polynomial function $f_{\varepsilon}$ that approximates $f$ within $\varepsilon/2$
$f_{\varepsilon}(\mathbf{x})=\sum_{\mathbf{j}\in M_{d}}\sum_{|\bm{\alpha}|\leq n-1}\underbrace{\frac{D^{\bm{\alpha}}f(\mathbf{c}_{\mathbf{j}})}{\bm{\alpha}!}}_{C_{\mathbf{j},\bm{\alpha}}}\psi_{\mathbf{j}}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\bm{\alpha}},\qquad\text{with}\ C_{\mathbf{j},\bm{\alpha}}\leq 1$
  • According to Theorem 3, we approximate each term $\psi_{\mathbf{j}}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\bm{\alpha}}=\prod_{i=1}^{d}(x_{i}-c_{j_{i}})^{\alpha_{i}}$, a product of $d+n$ factors
    $\mathbf{1}_{[\alpha_{j_{1}},\beta_{j_{1}}]}(x_{1})\times\mathbf{1}_{[\alpha_{j_{2}},\beta_{j_{2}}]}(x_{2})\times\cdots\times\mathbf{1}_{[\alpha_{j_{d}},\beta_{j_{d}}]}(x_{d})\times(x_{1}-c_{j_{1}})\times\cdots\times(x_{2}-c_{j_{2}})\times\cdots\times(x_{d}-c_{j_{d}})\times\cdots$
    by the network
    $\widetilde{G}_{\mathbf{j},\bm{\alpha}}(\mathbf{x})=\widetilde{M}_{k}\left(\phi_{\alpha_{j_{1}},\beta_{j_{1}},\tau,\kappa}(x_{1}),\widetilde{M}_{k}\left(\phi_{\alpha_{j_{2}},\beta_{j_{2}},\tau,\kappa}(x_{2}),\ldots,\widetilde{M}_{k}\left((x_{1}-c_{j_{1}}),(x_{1}-c_{j_{1}})\right)\right),\ldots\right)$
    We assume that the approximation error for a product is $\eta$ i.e. $|\widetilde{M}_{k}(x,y)-xy|\leq\eta$ (depth $O(1)$ and size $O(1)$).
    Let us analyze the propagation of approximation errors for $(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\bm{\alpha}}.$ Denote by $\gamma_{i},i\geq 2$ the approximation error at step of i1i-1 factors with $\gamma_{2}=0.$
    $\gamma_{i+1}=\left|\widetilde{M}_{k}\left(u_{i+1},\widetilde{\prod_{i=1}^{i}u_{i}}\right)-u_{i+1}\prod_{i=1}^{i}u_{i}\right|$ $\leq\left|\widetilde{M}_{k}\left(u_{i+1},\widetilde{\prod_{i=1}^{i}u_{i}}\right)-u_{i+1}\widetilde{\prod_{i=1}^{i}u_{i}}\right|+\left|u_{i+1}\widetilde{\prod_{i=1}^{i}u_{i}}-u_{i+1}\prod_{i=1}^{i}u_{i}\right|$
    $\leq\begin{cases}\eta+\gamma_{i},\ \text{If}\ \ |u_{i+1}|=|\phi_{\alpha_{j_{i+1}},\beta_{j_{i+1}},\tau,\kappa}(x_{i+1})|\leq 1\\ \eta+2\gamma_{i},\ \text{If}\ \ |u_{i+1}|=|x_{i+1}-c_{j_{i}}|\leq B\leq 2,\end{cases}$
    During successive approximation, the corresponding error evolves linearly:
    $\varepsilon_{\text{m}}=\eta\sum_{j=n}^{d+n-2}1^{j}+\eta\sum_{j=0}^{n-2}2^{j}\leq d\eta+\eta(2^{n-2}-1).$
    So the network $\widetilde{G}_{\mathbf{j},\bm{\alpha}}(\mathbf{x})$ approximates the term $\psi_{\mathbf{j}}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\bm{\alpha}}$ with error
    $\|\psi_{\mathbf{j}}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\bm{\alpha}}-\widetilde{G}_{\mathbf{j},\bm{\alpha}}(\mathbf{x})\|_{\infty}\leq\eta(2^{n-2}-1+d).$
Summing over all terms $(\mathbf{j},\bm{\alpha})$ (finite number), we obtain
$\sup_{\mathbf{x}\in[-B,B]^{d}}\big|\widetilde{f}_{\varepsilon}(\mathbf{x})-f_{\varepsilon}(\mathbf{x})\big|$ $\leq\big|\sum_{\mathbf{j},\bm{\alpha}}\left(C_{\mathbf{j},\bm{\alpha}}\ \widetilde{G}_{\mathbf{j},\bm{\alpha}}(\mathbf{x})-\psi_{\mathbf{j}}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\bm{\alpha}}\right)\big|$
$\leq\big|\sum_{\mathbf{j}:\mathbf{x}\in\text{supp}\psi_{\mathbf{j}}}\sum_{\bm{\alpha}}\left(C_{\mathbf{j},\bm{\alpha}}\ \widetilde{G}_{\mathbf{j},\bm{\alpha}}(\mathbf{x})-\psi_{\mathbf{j}}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\bm{\alpha}}\right)\big|$
$\leq 2^{d}\max_{\mathbf{j}:\mathbf{x}\in\text{supp}\psi_{\mathbf{j}}}\sum_{\bm{\alpha}}\left(C_{\mathbf{j},\bm{\alpha}}\ \big|\widetilde{G}_{\mathbf{j},\bm{\alpha}}(\mathbf{x})-\psi_{\mathbf{j}}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\bm{\alpha}}\big|\right)$
$\leq 2^{d}d^{n}\eta(2^{n-2}-1+d)$
Then, it suffices to choose $\eta$ such that
$\eta=\frac{\varepsilon}{2^{d+1}d^{n}(2^{n-2}-1+d)}.$
i.e.
$k\geq\max\left\{\left\lceil\frac{\ln C_{2}-\ln\eta}{2\ln\omega_{2}}\right\rceil,\left\lceil\frac{\ln 2B}{2\ln\omega_{2}}\right\lceil\right\}=t_{0}(n,d)(\ln 1/\varepsilon+t_{1}(n,d))$
to guarantee
$\sup_{\mathbf{x}\in[-B,B]^{d}}\big|\widetilde{f}_{\varepsilon}(\mathbf{x})-f_{\varepsilon}(\mathbf{x})\big|\leq\frac{\varepsilon}{2}.$
Thus, we show that the network $\widetilde{f}_{\varepsilon}$ approximates the function $f$ with error $\varepsilon$
$\sup_{\mathbf{x}\in[-B,B]^{d}}\big|\widetilde{f}_{\varepsilon}(\mathbf{x})-f(\mathbf{x})\big|$ $\leq\sup_{\mathbf{x}\in[-B,B]^{d}}\big|\widetilde{f}_{\varepsilon}(\mathbf{x})-f_{\varepsilon}(\mathbf{x})+f_{\varepsilon}(\mathbf{x})-f(\mathbf{x})\big|$
$\leq\sup_{\mathbf{x}\in[-B,B]^{d}}\big|\widetilde{f}_{\varepsilon}(\mathbf{x})-f_{\varepsilon}(\mathbf{x})\big|+\sup_{\mathbf{x}\in[-B,B]^{d}}\big|f_{\varepsilon}(\mathbf{x})-f(\mathbf{x})\big|$
$\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}$
$\leq\varepsilon.$
We construct the network by partitioning $B,B]^{d}$ into $M^{d}=O(\varepsilon^{-d/n})$ cubes through the choice $M=\lceil B(2^{d+1}/(n!\varepsilon))^{1/n}\rceil$, and locally approximate $f$ by its Taylor polynomial $T_{\mathbf{j}}$ of order n1n-1 at each center $\mathbf{c}_{\mathbf{j}}$, then define the global approximant $f_{\varepsilon}(\mathbf{x})=\sum_{\mathbf{j}}\psi_{\mathbf{j}}(\mathbf{x})T_{\mathbf{j}}(\mathbf{x})$ where $\psi_{\mathbf{j}}$ is the indicator function of cube $C_{\mathbf{j}}$, ensuring $\|f-f_{\varepsilon}\|_{\infty}\leq\varepsilon/2$. Each term $\psi_{\mathbf{j}}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{\mathbf{j}})^{\alpha}$ is a product of at most d+n1d+n-1 factors, approximated by a network $\widetilde{G}_{\mathbf{j},\alpha}$ using an approximate multiplication $\widetilde{M}_{k}$ of depth 1 and width 2, with per-multiplication error $\eta$ chosen such that $\eta(2^{d+n-1}-1)\leq\varepsilon/(2^{d+1}d^{n})$ by taking $k\sim\ln(1/\varepsilon)$. The $\mathcal{O}(\varepsilon^{-d/n})$ terms $\widetilde{G}_{\mathbf{j},\alpha}$ are computed in parallel over $\mathcal{O}(1)$ layers with width $\mathcal{O}(\varepsilon^{-d/n})$, then summed linearly via the addition of one linear layer, yielding a network $\widetilde{f}_{\varepsilon}$ with total depth $\mathcal{O}(1)+1=c(d,n,B)$ and size $\mathcal{O}(\varepsilon^{-d/n})$ satisfying $\|f-\widetilde{f}_{\varepsilon}\|_{\infty}\leq\varepsilon$. ∎
Appendix B Finite Difference Operator $\Delta$
The finite difference operator $\Delta$ is define in [jordan1965calculus].
Let $\Delta f(x)=f(x+1)-f(x)$. Noting that $\Delta=E-I$ where $Ef(x)=f(x+1)$ and $If(x)=f(x)$, the binomial theorem of Newton allows us to write
$\Delta^{m}=(E-I)^{m}=\sum_{j=0}^{m}\binom{m}{j}E^{j}(-I)^{m-j}$
Since $E^{j}f(0)=f(j)$, we finally obtain the formula:
$\Delta^{m}f(0)=\sum_{j=0}^{m}(-1)^{m-j}\binom{m}{j}f(j)$
Let us recall a useful property of finite differences of falling factorials:
$x^{\underline{k}}=x(x-1)\dots(x-k+1)$
of degree $k$, leading coefficient 1.
Lemme 2.
$\Delta x^{\underline{k}}=kx^{\underline{k-1}}$ and $\Delta^{m}x^{\underline{k}}=k(k-1)\dots(k-m+1)x^{\underline{k-m}}\quad\text{if }k\geq m$
Proof.
$\Delta x^{\underline{k}}=(x+1)^{\underline{k}}-x^{\underline{k}}$
Now $(x+1)^{\underline{k}}=(x+1)x(x-1)\dots(x-k+2)$ and $x^{\underline{k}}=x(x-1)\dots(x-k+1)$.
Factoring by $x(x-1)\dots(x-k+2)$, the difference is written
$\Delta x^{\underline{k}}=[(x+1)-(x-k+1)]\cdot x(x-1)\dots(x-k+2)=k\cdot x^{\underline{k-1}}$
We prove the second equality by induction. for $m=1,$ the second equality is true because it is equivalent to the first. Assume that
$\Delta^{m}x^{\underline{k}}=k(k-1)\dots(k-m+1)\,x^{\underline{k-m}},\quad k\geq m.$
is true. Then
$\Delta^{m+1}x^{\underline{k}}=\Delta\left[k(k-1)\dots(k-m+1)\,x^{\underline{k-m}}\right].$
The constant factor $k^{\underline{m}}:=k(k-1)\dots(k-m+1)$ is factored out of the operator $\Delta$:
$\Delta^{m+1}x^{\underline{k}}=k^{\underline{m}}\cdot\Delta x^{\underline{k-m}}.$
But
$\Delta x^{\underline{k-m}}=(k-m)\,x^{\underline{k-m-1}}.$
Therefore
$\Delta^{m+1}x^{\underline{k}}=k^{\underline{m}}\cdot(k-m)\,x^{\underline{k-m-1}}=k^{\underline{m+1}}\,x^{\underline{k-(m+1)}}.$
Appendix C An upper bound of $|\sigma^{(n)}(a)|$
The function $\sigma$ is analytic in $\mathbb{C}$, in particular in a disk $D(a,r)$,
$\sigma^{(n)}(a)=\frac{n!}{2\pi i}\int_{|x-a|=r}\frac{\sigma(x)}{(x-a)^{n+1}}dx.$
An upper bound gives
$|\sigma^{(n)}(a)|=\left|\frac{n!}{2\pi i}\int_{|x-a|=r}\frac{\sigma(x)}{(x-a)^{n+1}},dx\right|$ $\leq\frac{n!}{2\pi}\int_{|x-a|=r}\left|\frac{\sigma(x)}{(x-a)^{n+1}}\right||dx|$
$\leq\frac{n!}{2\pi r^{n+1}}\left(\max_{|x-a|=r}|\sigma(x)|\right)\int_{|x-a|=r}|dx|$
$\leq\frac{n!}{r^{n}}\left(\max_{|x-a|=r}|\sigma(x)|\right)$
The poles of $\sigma(x)$ are solutions of $1+e^{-x}=0$, i.e., $e^{-x}=-1$, so $x_{k}=i\pi+2ik\pi$, $k\in\mathbb{Z}$. The distance from $a\in\mathbb{R}$ to the nearest poles is $d_{a,\pi}=\sqrt{a^{2}+\pi^{2}}\geq\pi$. If $r<d_{a,\pi}$, then in the circle $\mathcal{C}_{r}=\{x\in\mathbb{C},|x-a|=r\}$, the function $x\mapsto|1+e^{-x}|$ is a continuous function (because there are no poles on the circle if $r<d_{a,\pi}$), so it attains a minimum $\min_{|x-a|=r}|1+e^{-x}|=\delta(a,r)>0$ on this compact set. Then for all $x$ on this circle,
$|1+e^{-x}|\geq\delta(a,r)\quad\Rightarrow\quad\frac{1}{|1+e^{-x}|}\leq\frac{1}{\delta(a,r)}.$
Consequently,
$|\sigma(x)|\leq\frac{1}{\delta(a,r)}\quad\text{for }x\ \text{such that}|x-a|=r,$
with $r<d_{a,\pi}.$ and it follows the upper bound
$|\sigma^{(n)}(a)|\leq C_{a}n!,$ (13)
where $C_{a}=\frac{1}{\delta(a,r)r^{n}}.$